### v2

# Additive Models Explained: A Computational Complexity Approach

Abstract
[…]
1 Introduction
Generalized additive models (GAMs) are a widely utilized family of models in ML (Hastie, 2017; Nelder and Wedderburn, 1972), valued commonly for their interpretability (Molnar, 2020; Caruana et al., 2015; Zhong et al., 2023; Bechler-Speicher et al., 2024; Chang et al., 2022; Liu et al., 2022; Bordt and von Luxburg, 2023; Enouen and Liu, 2025; Chen et al., 2020; McTavish et al., 2024; Mueller et al., 2024; Enouen and Liu, 2022; Zhuang et al., 2021) due to their additive structure. Beyond being considered interpretable, the very assumption of “additive interpretability” is a foundational one in the explainable AI community. For instance, many explanation techniques are designed to approximate additive behaviors (Ribeiro et al., 2016), and metrics such as infidelity (Yeh et al., 2019) assess explanation quality by measuring how well an additive approximation captures the effects of feature perturbations on the model’s output. Furthermore, interpretable training approaches often seek to promote near-additive behavior in models within specific domains (Alvarez Melis and Jaakkola, 2018).
Since GAMs are commonly regarded as an interpretable class of ML models, it may seem natural to expect that different types of explanations for their predictions can be computed efficiently, without encountering drastic computational intractability (e.g., NP-hardness, #\#P-hardness, etc.). This question also relates to a growing body of research in the ML community focused on understanding the computational complexity involved in generating different types of explanations for various model classes (Barceló et al., 2020; Wäldchen et al., 2021; Bassan et al., 2024). These efforts span both inherently complex models like neural networks (Adolfi et al., 2025) and tree ensembles (Audemard et al., 2022), where explanations are expected to be computationally hard to compute (Barceló et al., 2020; Adolfi et al., 2025), as well as traditionally “interpretable” models such as decision trees (Arenas et al., 2022) and monotonic classifiers (Marques-Silva et al., 2020), where computing explanations is expected to be more efficient (Barceló et al., 2020; Marques-Silva and Ignatiev, 2023; Bassan et al., 2024).
Our contributions. In this work, we investigate the computational complexity of generating various types of explanations for different classes of GAMs across a range of contexts. We also question the assumption that such explanations are inherently efficient to compute, given the perceived interpretability of GAMs. Through a comprehensive analysis spanning a wide array of settings, we uncover a surprisingly diverse landscape — revealing both tractable (efficient) and intractable (hard) cases. Our findings reveal that the complexity of generating explanations is heavily influenced by several factors — including some unexpected ones that were not known to impact complexity in previously studied ML models.
Our analysis is structured around the following three dimensions:
  1. 1.
    Component model type used within the GAM: Here, we investigate several popular component models that are typically used in GAMs and span the interpretability spectrum, including: (i) Smooth GAMs— i.e., the classical use of splines (Hastie, 2017); (ii) Neural Additive Models (NAMs)(Agarwal et al., 2021)— where each component is a neural network; and (iii) Explainable Boosting Machines (EBMs)(Caruana et al., 2015)— which use boosted tree ensembles as their components.
  2. 2.
    Type of explanation considered: We study the complexity of generating a diverse set of explanation types, including: (i) Minimum sufficient reasons— a common feature selection task that considers selecting the minimal subset of features that satisfy the common sufficiency criterion (as well as two additional relaxations of this task); (ii) Minimum contrastive explanations— minimal changes to the input that would alter the prediction; (iii) Shapley value attributions— assigning importance scores to features; and (iv) Feature redundancy identification— detecting redundant feature contributions.
  3. 3.
    Input domain setting: We study three general types of input domains: (i) Enumerable discrete domain— where each input variable takes values from a constant set; (ii) General discrete domain— where input values are discrete but not necessarily enumerable; (iii) Continuous domain— where inputs take real-valued features.
We prove complexity results for generating explanations across all configurations of these dimensions — i.e., across every combination of component type $\times$ explanation type $\times$ input domain (see Table 1 for a summary of the results). This uncovers a diverse range of results — including (i) efficient polynomial-time algorithms, (ii) intractabilityresults (NP-hard, #P-hard, etc.), and (iii) pseudo-polynomial timeresults — cases where the problem is generally intractable but becomes efficiently solvable when the GAM’s weight coefficients are encoded in unary rather than binary. Intuitively, pseudo-polynomial tractability suggests that although generating explanations for a GAM may be intractable, reducing the precision of its coefficient weights can render the explanation task efficient.
Key Insights. Beyond the direct contributions that our results offer of both efficient algorithms and intractability outcomes, they also uncover several surprising insights into the computational nature of generating explanations for GAMs, as highlighted below:
  • The complexity of computing explanations for GAMs depends heavily on the input domain, unlike other ML models where a variation based on the input domain is not observed. We show that for most explanation types we studied — like sufficient explanations, contrastive explanations, and Shapley values — computing explanations becomes exponentially harder in continuous and discrete settings compared to enumerable discrete ones. An interesting exception is the feature redundancy explanation, which is actually exponentially easier in the continuous case. This significant sensitivity to the input domain is surprising since it appears to be unique to additive models, as other ML models (e.g., decision trees, tree ensembles, neural networks) do not exhibit such diversity in complexity across input domains.
  • The complexity of obtaining explanations for GAMs largely depends on their component models, but this effect interestingly appears only in certain input domains. We show that for all our explanation types, there are exponential differences in the complexity of generating explanations for GAMs, depending on the underlying component models (e.g., GAMs with splines are significantly easier to obtain explanations for than NAMs and EBMs). Interestingly, however, we uncover that this complexity distinction arises only in the continuous and general discrete input domains, but not in the enumerable discrete domain.
  • An intriguing complexity separation shows that computing explanations for GAMs used for classification is strictly harder than for regression — but only for SHAP, not other explanations. We demonstrate this by proving polynomial-time results for the regression setting, in contrast to the classification setting, which is #P-Hard — highlighting a notable complexity separation.
  • Some non-additive models (e.g., neural networks) are harder to obtain explanations over compared to their additive counterparts (e.g., NAMs), but this depends critically on the explanation type and input domain. We show that models typically hard to interpret — such as tree ensembles and neural networks, where computing explanations is at least NP-Hard — become tractable (e.g., solvable in polynomial time, or belonging to lower complexity classes) in their additive forms (e.g., NAMs, EBMs). Interestingly, however, this phenomenon does not always occur and is highly sensitive to the type of explanation and the input domain.
Due to space limitations, we provide only a brief outline of the proofs for our claims in Section 6, with the full proofs presented in the appendix.
2 Preliminaries
Complexity Classes. This paper assumes familiarity with basic complexity classes such as polynomial-time (PTIME), and non-deterministic polynomial-time (NP, coNP). We also discuss the class $\Sigma_{2}^{P}$, which includes problems solvable in NP with access to a coNP oracle and $\Delta^{P}_{2}$, which includes problems solvable in PTIME with access to an NP oracle. Clearly, NP and coNP are contained in $\Delta^{P}_{2}$ and $\Sigma_{2}^{P}$, and it is also widely believed that NP, coNP $\subsetneq\Delta^{P}_{2}\subsetneq$ $\Sigma_{2}^{P}$ Arora and Barak (2009). Additionally, we cover #P, the class of functions that count the accepting paths of polynomial-time nondeterministic Turing machines. Although not directly comparable since they focus on counting rather than decision problems, it is widely believed that the #P class is strictly “harder” than $\Sigma_{2}^{P}$, which can be denoted as $\Sigma_{2}^{P}$ $\subsetneq$#P Arora and Barak (2009).
Setting. We consider a set of input features indexed by $\{1,\ldots,n\}$, where $n\in\mathbb{N}$, and denote an input instance as $\textbf{x}:=(\textbf{x}_{1},\ldots,\textbf{x}_{n})$. Each feature $i$ has a domain $\mathcal{X}_{i}$, such that the input space is defined as $\mathbb{F}:=\mathcal{X}_{1}\times\mathcal{X}_{2}\times\ldots\times\mathcal{X}_{n}$. The predictive model $f$ maps inputs from $\mathbb{F}$ either to a regression output, $f:\mathbb{F}\to\mathbb{R}^{d}$ with dimension $d\in\mathbb{N}$, or to a classification label, $f:\mathbb{F}\to[c]$, where $c\in\mathbb{N}$ is the number of classes. For clarity in our proofs, we restrict $f$ to have a one-dimensional output — i.e., $f:\mathbb{F}\to\mathbb{R}$ in the regression case and $f:\mathbb{F}\to\{0,1\}$ in the classification case — though we emphasize that our results extend to multi-dimensional outputs (see Appendix D for details). The majority of explanations we study are local, aiming to interpret the model’s prediction at a specific instance $\textbf{x}\in\mathbb{F}$ — that is, to explain why $f(\textbf{x})$ was predicted. However, we also consider some global explanations, which aim to interpret the overall behavior of $f$, independent of any particular input.
Generalized Additive Models (GAMs). A GAM $f$ is defined as an additive model consisting of $k$ component models $f_{1},f_{2},\ldots,f_{k}$. Formally, in the regression setting, we define:
$f(\textbf{x}):=\beta_{0}+\beta_{1}f_{1}(\textbf{x}_{1})+\beta_{2}f_{2}(\textbf{x}_{2})+\ldots\beta_{k}f_{k}(\textbf{x}_{k})$ (1)
where $\beta_{0},\beta_{1},\ldots,\beta_{k}\in\mathbb{Q}$ are the intercept terms and $f(\textbf{x})\in\mathbb{R}$ is the regressor predictive value. For the scenario where $f(\textbf{x})$ is a classification model, we assume a step function over the additive prediction, i.e., $f(\textbf{x}):=\text{step}(\beta_{0}+\beta_{1}f_{1}(\textbf{x}_{1})+\beta_{2}f_{2}(\textbf{x}_{2})+\ldots\beta_{k}f_{k}(\textbf{x}_{k}))$ where we define the step function as $\text{step}(\textbf{z})=1$ if $\textbf{z}\geq 0$, and $\text{step}(\textbf{z})=0$ otherwise.
3 Dimensions of the Complexity Analysis
As previously mentioned, our detailed complexity analysis spans three distinct dimensions: (i) the input domain, (ii) the types of component models utilized within GAMs, and (iii) the explanation types . In this section, we elaborate on each of these dimensions.
3.1 Dimension 1: The input domain
We study the following cases for the domain $\mathcal{X}_{i}$ of feature $i\in\{1,\ldots n\}$: (i) when the domain is enumerable discrete, meaning each value $\textbf{x}_{i}$ is selected from a given set $\mathcal{K}\subseteq\mathbb{Q}$ of a constant number $|\mathcal{K}|$ of possible assignments. Intuitively, these are scenarios where each input is limited to a predefined set of values, which is often relevant for cases such as tabular data or certain language tasks; (ii) the discrete setting, or the general discrete setting, where $\mathcal{X}_{i}$ represents any discrete input within a range defined by maximum and minimum values expressed in binary with $q$ bits. This intuitively allows a broader range of discrete assignments for each input, while still operating within a discrete domain. (iii) The continuous setting, where $\mathcal{X}_{i}\subseteq\mathbb{R}$, which ensures guarantees across the entire infinite spectrum of the specified domain . Both discrete and continuous domains are widely used across various ML tasks, including those in vision and language tasks (Elkabetz and Cohen, 2021; Hong et al., 2022; Cartuyvels et al., 2021). We note that general discrete settings naturally encompass enumerable ones and are thus always at least as computationally challenging, whereas continuous settings do not enable a direct comparison to either and can make problems either harder or easier to solve (Karmarkar, 1984; Vavasis, 1991; Arora and Barak, 2009).
3.2 Dimension 2: The component model type
We consider a diverse range of component models that can be integrated into a GAM, spanning various levels of interpretability and reflecting commonly used approaches in the literature: (i) The classic spline-based models (Hastie, 2017), where each component function is piecewise polynomial. In practice, these polynomials are typically restricted to a degree of at most three, yielding the widely used cubic splines (Hastie, 2017), which will be our primary focus (though our results can often be extended to higher-order splines — see Appendix B). To distinguish these GAMs from the non-smooth variants, we refer to GAMs with spline-based components as Smooth GAMs. (ii) Neural Additive Models (NAMs)(Agarwal et al., 2021; Bouchiat et al., 2023; Xu et al., 2023; Jiao et al., 2024; Thielmann et al., 2024), in which each component is a neural network with ReLU activations. (iii) Explainable Boosting Machines (EBMs)(Caruana et al., 2015; Wei et al., 2022; Nori et al., 2021), where each component function is a boosted tree ensemble. We present the complete formal definitions of all model types in Appendix B.
3.3 Dimension 3: The explanation type
To explore various dimensions of GAM interpretability, we examine a broad range of widely used explanation types drawn from the existing literature. Building upon prior computational complexity frameworks for evaluating explanations in ML models (Barceló et al., 2020; Arenas et al., 2021a; Bassan et al., 2024), we formalize each type of explanation as an explainability query. Such a query takes a model $f$ and x as inputs and aims to address specific questions, offering a meaningful interpretation of the prediction $f(\textbf{x})$.
Sufficient reason Feature Selection. We consider the common sufficiency criterion for feature selection, used in popular explainability methods (Ribeiro et al., 2018; Carter et al., 2019; Ignatiev et al., 2019b). A sufficient reason is a subset of input features, $S\subseteq[n]$, such that fixing the features in $S$ to their values in $\textbf{x}\in\mathbb{F}$ guarantees the prediction remains $f(\textbf{x})$, regardless of the assignment to $\overline{S}$. We write $(\mathbf{x}_{S};\mathbf{z}_{\bar{S}})$ for an input where $S$ takes values from x and $\overline{S}$ from z. Formally, $S$ is a sufficient reason for $\langle f,\textbf{x}\rangle$ iff for all $\textbf{z}\in\mathbb{F}$: $f(\textbf{x}_{S};\textbf{z}_{\bar{S}})=f(\textbf{x})$.
A common assumption in the literature is that smaller sufficient reasons (those with smaller $|S|$) are more useful (Ribeiro et al., 2018; Carter et al., 2019; Ignatiev et al., 2019b). This motivates the search for cardinally minimal sufficient reasons, also called minimum sufficient reasons, and leads to our first explainability query:
MSR (Minimum Sufficient Reason): Input: Model $f$, input x, and $d\in\mathbb{N}$ Output: Yes if there exists some $S\subseteq[n]$ such that $S$ is a sufficient reason with respect to $\langle f,\textbf{x}\rangle$ and $|S|\leq d$, and No otherwise.
To deepen our understanding of the complexity of sufficient reasons, we also examine two common related explainability queries that refine the MSR query (Barceló et al., 2020; Bassan et al., 2024): (i) Check-Sufficient-Reason(CSR), which checks if a given subset $S$ is a sufficient reason; (ii) Count-Completions(CC), a generalized version of CSR that measures the fraction of completions preserving the prediction, capturing the probability of maintaining the classification. Full formalizations appear in Appendix C.
Contrastive Explanations. An alternative way to interpret models is by identifying subsets of features that, when changed, could alter the model’s prediction (Dhurandhar et al., 2018; Guidotti, 2022). We define s subset $S\subseteq[n]$ as contrastive if modifying its values can change the classification $f(\textbf{x})$, i.e., there exists $\textbf{z}\in\mathbb{F}$ such that $f(\textbf{x}_{\bar{S}};\textbf{z}_{S})\neq f(\textbf{x})$. As with sufficient reasons, smaller contrastive subsets are typically assumed to be more meaningful (Dhurandhar et al., 2018; Guidotti, 2022; Ignatiev et al., 2020; Mothilal et al., 2020), motivating a focus on cardinally-minimal contrastive reasons:
MCR (Minimum Change Required): Input: Model $f$, input x, and $d\in\mathbb{N}$. Output: Yes, if there exists some contrastive reason $S$ such that $|S|\leq d$ for $f(\textbf{x})$, and No otherwise.
Shapley Values. In the additive attribution setting, each feature $i\in[n]$ is assigned an importance weight $\phi_{i}$. A common method for assigning these weights is the Shapley value attribution index (Lundberg and Lee, 2017):
$\phi_{i}(f,\textbf{x}):=\sum_{S\subseteq[n]\setminus\{i\}}\frac{|S|!(n-|S|-1)!}{n!}(v(S\cup\{i\})-v(S))$ (2)
We use the standard conditional expectation value function $v(S):=\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f(\textbf{z})\mid\textbf{z}_{S}=\textbf{x}_{S}]$ (Sundararajan and Najmi, 2020; Lundberg and Lee, 2017). Moreover, as is standard in both complexity analyses (Arenas et al., 2023; Van den Broeck et al., 2022) and practical SHAP methods like KernelSHAP (Lundberg and Lee, 2017), we assume feature independence. See Appendix C for a full formalization.
SHAP (Shapley Additive Explanation): Input: Model $f$, input x, and $i\in[n]$ Output: The shapley value $\phi_{i}(f,\textbf{x})$.
Feature redundancy. For our final explanation form, we study the complexity of determining whether a feature $i$ is redundant with respect to a model $f$, following the notion of redundancy explored in prior work (Arenas et al., 2021a; Bassan et al., 2024; Huang et al., 2023; Darwiche and Hirth, 2020). Unlike the previous local explanation forms, this one is global — it checks if $i$ is redundant for all $\textbf{x}\in\mathbb{F}$. Formally, $i$ is redundant if for all $\textbf{x},\textbf{z}\in\mathbb{F}$, we have $f(\textbf{x})=f(\textbf{x}_{[n]\setminus\{i\}};\textbf{z}_{\{i\}})$.
FR (Feature Redundancy): Input: Model $f$, and integer $i$. Output: Yes, if $i$ is redundant with respect to $f$, and No otherwise.
4 Main Complexity Results
We examine the complexity across all analyzed settings — spanning input domains (enumerable discrete, discrete, continuous), model types (NAMs, EBMs, Smooth GAMs), and explanation forms (CSR, MSR, MCR, FR, CC, SHAP). We distinguish between SHAP for regression and classification (SHAP-R vs. SHAP-C), a complexity split, which, as we explain in detail later, is unique to SHAP. For each setting, we provide a novel complexity proof, beginning with a summary in Table 1.
Table 1: A summary of complexity results across all component-model types, input domains, and explanations. Bold font emphasizes the most tractable complexity classes for each explanation.
Input Space Component Type CSR, MSR MCR FR CC, SHAP-C SHAP-R
Enumerable Discrete Any PTIME PTIME coNP-C #\#P-C (Pseudo-P) PTIME
Discrete Smooth GAMs PTIME PTIME coNP-C #\#P-C PTIME
NAMs, EBMs coNP-H NP-C coNP-C #\#P-C #\#P-C
Continuous Smooth GAMs PTIME PTIME PTIME #\#P-C PTIME
NAMs, EBMs coNP-H NP-C coNP-C #\#P-C #\#P-C
The results in Table 1 identify when explanations can be computed efficiently (e.g., in polynomial or pseudo-polynomial time) and when they are computationally intractable (e.g., NP-Hard). A central finding is the significant variation in complexity, influenced by factors such as the input domain, the type of component model, and whether the task is regression or classification. In Section 5, we analyze this diverse landscape in depth and compare these complexity results to existing complexity results for non-additive models like neural networks and tree ensembles. In Section 6, we will outline the proofs underlying these complexity results.
5 A Computational Interpretability Hierarchy for GAMs
Our results uncover a rich and varied landscape of complexity outcomes across different settings, which we analyze in this section. To present these complexity distinctions more elegantly, we adopt the notation from (Barceló et al., 2020), indicating when one model is strictly more computationally interpretable (c-interpretable) than another — that is, when it belongs to a strictly lower complexity class.
Definition 1.
Let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ be two classes of models and let $Q$ be an explainability query for which $Q(\mathcal{C}_{1})$ is in complexity class $\mathcal{K}_{1}$ and $Q(\mathcal{C}_{2})$ is in complexity class $\mathcal{K}_{2}$. We say that $\mathcal{C}_{1}$ is strictly more c-interpretable than $\mathcal{C}_{2}$ with respect to $Q$ iff $Q(\mathcal{C}_{2})$ is hard for the complexity class $\mathcal{K}_{2}$ and $\mathcal{K}_{1}\subsetneq\mathcal{K}_{2}$.
5.1 The Sensitivity of Complexity to the Input Domain, in Contrast to Other ML Models
Our results reveal significant exponential complexity distinctions in explaining GAMs based on the input domain — a surprising finding, as such distinctions do not appear in other popular ML models (e.g., decision trees, tree ensembles, neural networks; see Appendix D for details). While explanations over enumerable discrete domains are often computable in polynomial time, they become intractable in general discrete or continuous settings. The only exception is the feature redundancy query.
Theorem 1.
GAMs over enumerable discrete settings are strictly more c-interpretable than GAMs over general discrete or continuous settings with respect to CSR, MCR, MSR, CC and SHAP.
The intuition is that in an enumerable discrete setting, the input space can be explicitly traversed to evaluate each component $f_{i}$, allowing computations such as minimum or maximum component output computation — key for sufficient or contrastive reason queries — as well as expected value computation, which is essential for SHAP. However, in the general discrete or continuous setting, such iteration is infeasible, making these computations intractable.
However, interestingly, we discover that for one of the queries we analyzed — the Feature Redundancy (FR) query, which seeks to determine whether a feature does not contribute to a model’s prediction — this type of explanation is actually strictly easier to obtain in the continuous setting compared to the discrete or even the enumerable discrete settings:
Theorem 2.
GAMs over continuous input settings are strictly more c-interpretable than GAMs over enumerable discrete or general discrete input settings with respect to FR.
We highlight that these complexity gaps, tied to input domains, are unique to additive models. For other model types — decision trees, neural networks, and tree ensembles — no such gaps appear (see Appendix D for an elaborate discussion):
Observation 1.
While there exist strict complexity gaps for the CSR, MCR, CC, SHAP, and FR queries between the different input setting configurations for GAMs, such complexity gaps do not hold for other ML models such as decision trees, tree ensembles, and neural networks.
5.2 The Significance of the Component Models on Complexity Depends on the Input Domain
We show that while computing some explanations for certain GAMs with more interpretable component models (such as splines) can be done efficiently, these same tasks may sometimes become intractable for GAMs consisting of uninterpretable component models (e.g., neural networks or boosted trees). We show that this phenomenon holds only in the continuous and discrete input spaces:
Theorem 3.
GAMs over continuous or discrete inputs composed of splines (Smooth GAMs) are strictly more c-interpretable than GAMs with neural networks or tree ensembles (NAMs, EBMs).
However, we surprisingly demonstrate that this distinction between model components does not always hold and is dependent on the input domain. Specifically, we show that this gap emerges only in the continuous and discrete input settings but not in the enumerable discrete setting. The intuition behind this result is that the enumerability of the input space mitigates the complexity originating from the component model itself, unlike in the discrete or continuous domains:
Observation 2.
While there exist strict complexity distinctions between component models when computing explanations for GAMs in the discrete and continuous input domain, such complexity distinctions do not hold under the enumerable discrete input domain.
We also note that the result shown in this subsection, together with our findings in Subsection 5.1 on input domain sensitivity, offers a complementary perspective to empirical work such as (Chang et al., 2021). These studies demonstrate that both the choice of component model and the form of data representation (i.e., input representation) substantially influence the learned functional structure of GAMs. Since differing functional forms naturally correspond to different levels of computational complexity, this aligns with our observation that the complexity of generating explanations is highly sensitive to these modeling choices.
5.3 Complexity Distinctions Between Regression and Classification, but Only for SHAP
In the context of SHAP explanations, we reveal an intriguing complexity distinction between classification and regression GAMs. Specifically, we show that while SHAP explanations for regression tasks on Smooth GAMs (or general GAMs with enumerable discrete inputs) can be computed in polynomial time, the same problem becomes computationally intractable (#P-hard) for classification tasks. The intuition underlying this distinction is that the classification scenario introduces an additional “step” function, whose non-linearity violates the linearity axiom essential for establishing tractability in the regression setting. This result motivates the following:
Theorem 4.
Regression Smooth GAMs and regression GAMs with enumerable inputs are strictly more c-interpretable than classification Smooth GAMs and classification GAMs with enumerable inputs with respect to SHAP.
We note that while SHAP allows for a direct comparison between regression and classification tasks, adapting our other explanation types — originally defined for classification — to regression requires redefining them (as in (Wu et al., 2023; Izza et al., 2024)) to assess whether values lie within a specified $\delta$-range. From a complexity standpoint, these behave similarly to classification cases and do not produce the same distinctions observed with SHAP. We elaborate on this in Appendix D.
5.4 Additive vs. Non-Additive Complexity Depends on the Input Domain and Explanation
As noted in the introduction, the ML community often assumes that increasing a model’s additivity improves interpretability. For instance, NAMs (Agarwal et al., 2021; Radenovic et al., 2022) and EBMs (Caruana et al., 2015) are generally seen as more interpretable than standard neural networks and boosted trees, respectively. It thus seems natural to expect that explanations for additive models (e.g., NAMs) are easier to compute than their non-additive counterparts. Our findings support this, showing a clear complexity gap between black-box models and their additive counterparts. However, importantly, this gap is not universal — it strongly depends on the explanation type and the input domain. Table 2 summarizes prior complexity results for neural networks and tree ensembles, alongside our new results for NAMs and EBMs.
Table 2: Complexity distinctions in explaining non-additive black-box models — tree ensembles (TEs) or neural networks (NNs), compared to their additive counterparts — neural additive models (NAMs) and explainable boosting machines (EBMs). We emphasize in bold font cases where a strict complexity separation exists between an additive and a non-additive configuration.
Enumerable Discrete Discrete, Continuous
EBMs, NAMs TEs, NNs EBMs, NAMs TEs, NNs
CSR PTIME coNP-C coNP-C coNP-C
MSR PTIME $\mathbf{\Sigma^{P}_{2}}$-C coNP-H, in $\Delta^{P}_{2}$ $\mathbf{\Sigma^{P}_{2}}$-C
MCR PTIME NP-C NP-C NP-C
FR coNP-C coNP-C coNP-C coNP-C
CC, SHAP-C #P-C (Pseudo-P) #P-C (No Pseudo-P) #\#P-C #\#P-C
SHAP-R PTIME #P-C #\#P-C #\#P-C
Particularly, we show that in the enumerable discrete input domain, there exist many complexity gaps between the non-additive models — neural networks and tree ensembles, and their additive counterparts (NAMs and EBMs). However, in the discrete and continuous input settings, while such a complexity gap is apparent in the minimum-sufficient-reason (MSR) query, it is not apparent in the other queries. These observations lead us to the following corollary:
Theorem 5.
NAMs and EBMs over enumerable discrete input settings are strictly more c-interpretable than neural networks and tree ensembles with respect to CSR, MSR, MCR, and SHAP for regression. However, NAMs and EBMs over general discrete or continuous input settings are strictly more c-interpretable than tree ensembles and neural networks with respect to MSR.
6 Proof Outline of Complexity Results
We begin by summarizing the tractable results — explanations that can be computed in polynomial time — in Subsection 6.1. Next, in Subsection 6.2, we discuss a relaxed setting where tractability is preserved under unary-encoded weights, referred to as pseudo-polynomial time. Finally, Subsection 6.3 covers the intractable cases, such as NP-hardness and #\#P-hardness results.
6.1 Polynomial-Time Explanation Computations
Sufficient and contrastive reasons. We begin by analyzing the polynomial-time complexity of computing sufficient and contrastive reason queries (CSR, MSR, MCR). A key observation is that by identifying the minimal and maximal values each component $\beta_{i}\cdot f_{i}(\textbf{x}_{i})$ can take, we can rank features by their “importance”, measured by how much altering each feature deviates the prediction. A greedy algorithm, akin to those in (Barceló et al., 2020; Marques-Silva et al., 2020), then selects features by importance until a minimal sufficient or contrastive reason is found. CSR is simpler, requiring only a sufficiency check for a given subset. Thus, the main computational challenge lies in efficiently computing the min/max values of each $f_{i}(\textbf{x}_{i})$. For enumerable discrete inputs, this can be done by enumeration; for smooth GAMs, these values can be computed analytically, even with general discrete or continuous inputs.
Proposition 1.
Given any GAM over an enumerable discrete domain, or a Smooth GAM over a discrete or continuous domain, the CSR, MCR, and MSR queries are solvable in polynomial time.
Feature redundancy for Smooth GAMs in continuous settings. Unlike the discrete or enumerable discrete settings, where identifying strict feature redundancies is intractable (see Subsection 6.3), in continuous domains we show that a feature is strictly redundant if and only if its component $\beta_{i}\cdot f_{i}(\textbf{x}_{i})$ is identically zero — i.e., either $\beta_{i}=0$ or $f_{i}(\textbf{x}_{i})=0$ for all $\textbf{x}_{i}$. For Smooth GAMs, this can be checked efficiently in polynomial time. This leads to the following proposition:
Proposition 2.
Given a Smooth GAM $f$ over a continuous input space, the FR query can be obtained in polynomial time.
SHAP explanations in regression settings. When computing SHAP, polynomial-time computation is feasible only for GAMs in regression tasks (classification intractability is discussed in Subsection 6.3). Leveraging the linearity axiom of Shapley values (Lundberg and Lee, 2017), we reduce the computation to evaluating the computations of $\beta_{i}\cdot\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f_{i}(\textbf{x}_{i})]$. When the model is either discretely enumerable or supports efficient integration (e.g., in Smooth GAMs), these expectations can be computed in polynomial time.
Proposition 3.
Given any regression GAM with enumerable discrete inputs, or given a regression Smooth GAM with discrete or continuous inputs, then the SHAP query is solvable in polynomial time.
6.2 Pseudo Polynomial-Time Explanation Computations
Table 1 shows that both CC and SHAP-C queries are intractable. However, if the GAM’s internal encoding and input is given in unary (rather than binary) and the GAM has a polynomially bounded output, they become computable in polynomial time. This pseudo-polynomial setting enables efficient computation when the encoding precision is bounded, using dynamic programming similar to (Barceló et al., 2020), and detailed in our proofs. It applies to enumerable discrete domains (all component types), where enumeration makes $\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f(\textbf{x}_{i})]$ tractable.
Proposition 4.
Given any unary-encoded GAM over a discrete, enumerable input domain with polynomially bounded output, the CC and SHAP-C queries admit a pseudo-polynomial-time algorithm.
6.3 Intractable Explanation Computations
Sufficient and contrastive reasons. As shown in Subsection 6.1, finding a cardinally minimal sufficient or contrastive reason in GAMs requires computing the minimal and maximal values of each $f_{i}(\textbf{x}_{i})$ component. We show this is intractable for neural networks and tree ensembles, as their exponential search space makes the problem computationally hard. The main difficulty is proving hardness even for single input-output instances, where lower expressivity demands more intricate constructions. We overcome this by reducing the instance to an equivalent $n$-dimensional input, mapping continuous or general discrete regions to their relevant counterparts, yielding the following:
Proposition 5.
Let there be a NAM or an EBM over a discrete or continuous input space, then solving the CSR query is coNP-Complete, the MSR query is coNP-Hard, and the MCR query is NP-Complete.
Feature redundancy. In contrast to the continuous case, where redundancy reduces to checking if $\beta_{i}\cdot f_{i}(\textbf{x}_{i})=0$, in discrete settings it requires verifying whether this value falls below the decision boundary — a task that is often computationally intractable:
Proposition 6.
Given a NAM, a Smooth GAM or an EBM under an enumerable discrete or general discrete setting, then the FR query is coNP-Complete.
Intractability of SHAP and count completions. Finally, we show that computing the SHAP and CC queries is generally #\#P-Hard by reducing to the model counting problem (Van den Broeck et al., 2022), implying intractability for EBMs and NAMs. While other settings are also intractable (see Subsection 6.2), their complexity can be mitigated using pseudo-polynomial algorithms:
Proposition 7.
Given NAMs and EBMs over discrete or continuous settings, computing SHAP for regression tasks is #\#P-Complete. Moreover, given any GAM over an enumerable discrete, discrete or continuous setting — computing the CC and SHAP queries for classification are #\#P-Complete.
7 Practical Implications
Although our work is primarily theoretical, it provides several insights with practical relevance for practitioners working with GAMs. First, we identify a broad class of efficient polynomial-time algorithms for computing explanations with different guarantees across diverse GAM settings. In fact, all of the tractable algorithms presented in this work (included in the appendix) are not only polynomial-time, but even linear in the size of the GAM, underscoring their practical applicability.
Second, we show that techniques commonly used in machine learning practice can also make explanation methods computationally feasible. In particular, we identify two strategies that can transform otherwise hard explanation problems into tractable ones:
  1. 1.
    Input domain transformations. We find that the computational complexity of generating explanations is highly sensitive to the structure of the input domain. Small changes in the domain can shift problems from intractable to tractable. This insight motivates the use of input transformations, such as discretization, a standard step in many GAM pipelines (Caruana et al., 2015; Nori et al., 2021), to reshape the input space in ways that make explanation computation efficient.
  2. 2.
    Quantization. Our pseudo-polynomial algorithms show that explanation problems that are intractable in full generality become tractable when the GAM encoding is quantized. This suggests a promising practical approach: by quantizing model coefficients, one can enable efficient computation of explanations that would otherwise be infeasible.
8 Related Work
Our work builds upon prior research exploring the computational complexity of obtaining various types of explanations for ML models (Barceló et al., 2020; Arenas et al., 2022; Blanc et al., 2021; Adolfi et al., 2025; Blanc et al., 2022; Huang et al., 2023; Barceló et al., 2025; Bhattacharjee and Luxburg, 2024; Kozachinskiy, 2023; Marzouk and De La Higuera, 2024; Laber, 2024; Bounia, 2024; Xue et al., 2023; Jin et al., 2025). This area is closely tied to the subfield of interest known as formal explainable AI (Marques-Silva and Ignatiev, 2022; Bassan and Katz, 2023) which focuses on explanations with mathematical guarantees (Ignatiev et al., 2019b; Wu et al., 2023; Izza et al., 2021; Audemard et al., 2020; Darwiche and Hirth, 2020; Bassan et al., 2025a; Izza et al., 2023b; Bassan et al., 2023; Bounia, 2025), where analyzing the complexity of producing explanations with such guarantees plays a central role (Izza et al., 2023a; Ordyniak et al., 2023; Audemard et al., 2021; Cooper and Marques-Silva, 2021; Arenas et al., 2022; Ordyniak et al., 2024; Van den Broeck et al., 2022).
Previous work has mainly examined the complexity of generating explanations for specific ML models, including black-box models like neural networks (Adolfi et al., 2024; Barceló et al., 2020) and interpretable models like decision trees (Arenas et al., 2022; Bounia and Koriche, 2023; Huang et al., 2021). Closer to our work are studies on linear models (Barceló et al., 2020; Subercaseaux et al., 2025; Marques-Silva et al., 2020) — a simplified form of GAMs — typically under binary input-output settings. In contrast, we offer a comprehensive analysis of a broad class of GAMs, covering diverse component models, explanation types, input assumptions, and approximation results. We elaborate further on related works in Appendix A.
Lastly, we note that some terms in our work have sometimes been referred to differently in prior literature. For instance, while the term sufficient reasons is widely used (Barceló et al., 2020; Bassan et al., 2025d; Darwiche and Hirth, 2020), it is also referred to as abductive explanations (Ignatiev et al., 2019a). Moreover, subset-minimal sufficient reasons are closely related — though not identical — to prime implicants in Boolean formulas (Darwiche and Marquis, 2002; Darwiche and Hirth, 2020; Shih et al., 2018).Similarly, the CC query corresponds to the $\delta$-relevant set Wäldchen et al. (2021); Izza et al. (2021); Subercaseaux et al. (2025), which checks if the completion count exceeds a given threshold $\delta$.
9 Limitations and Future Work
As in prior work on the computational complexity of obtaining explanations, our analysis focuses on specific explanation definitions and component model types, though many other settings could certainly be explored. Nevertheless, we believe that our work provides a broad perspective across a wide array of popular explanation methods, component types, and input settings, highlighting various novel aspects of the explainability landscape for GAMs and establishing a strong foundation for exploring additional settings in future research.
More concretely, our results open the door to a wide range of future practical implementations of our algorithms (see Section 9). From a theoretical perspective, our analysis can be extended in several directions. A first promising avenue is to better understand how high-order interactions, which are widely used in GAMs (Chang et al., 2022; Lou et al., 2013; Kim et al., 2022; Enouen and Liu, 2025), affect the complexity of generating explanations, and how this varies across explanation types and component models. Another interesting direction is to study how structural assumptions over the GAM influence complexity. For instance, one may consider the effect of concurvity (Kovács, 2024; Siems et al., 2023). In the extreme case of high concurvity, where smooth components are strongly correlated and behave almost like a single spline, the GAM may resemble a simpler function class, potentially making some explanation problems easier. Conversely, in the case of low concurvity, especially for explanation types that rely on interactions, the high separability between components may also offer computational advantages. Finally, for intractable settings, one can attempt to leverage automated reasoning tools, such as neural network verifiers (Wu et al., 2024; Wang et al., 2021b) or MILP/MaxSAT solvers for tree ensembles (Ignatiev et al., 2022; Chen et al., 2019) to obtain explanations. Moreover, while we propose two central strategies for circumventing intractability, input transformations and quantization (see Section 9), exploring additional circumvention approaches such as approximation guarantees, probabilistic or statistical relaxations, and PAC-based guarantees remains an important direction for future research.
10 Conclusion
We present a theoretical framework for assessing the computational complexity involved in obtaining various forms of explanations for different types of GAMs. Our work uncovers a nuanced spectrum of results, demonstrating that the complexity of obtaining explanations for GAMs is significantly influenced by (i) the input space domain (enumerable discrete, discrete, and continuous), (ii) the type of explanation (sufficient, contrastive, shapley value explanations, etc.), (iii) the underlying component model (e.g., neural networks, boosted trees, splines, etc.), (iv) and the distinction between classification and regression settings. Our findings reveal a broad range of unexpected and substantial complexity variations, driven by surprising factors — such as the input domain — which are not apparent in other ML models. We believe that our work lays the foundation for a wide range of implementations by enabling the development of tractable algorithms for computing explanations for GAMs in diverse settings. At the same time, it advances the theoretical understanding of these models by identifying when such explanations are computationally feasible and when they are not.
Acknowledgments
This work was partially funded by the European Union (ERC, VeriDeL, 101112713). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This research was additionally supported by a grant from the Israeli Science Foundation (grant number 558/24).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction of this work directly reflect the proofs we establish and their implications.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We provide a dedicated section outlining the limitations.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Due to space limitations, we present only a proof outline in Section 6, with all full and detailed proofs deferred to the appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [N/A] .
  19. Justification: This is a theoretical work; thus, experimental reproducibility is not applicable.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [N/A] .
  24. Justification: This is a theoretical work; therefore, providing open access to data and code is not relevant.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [N/A] .
  29. Justification: This is a theoretical work; therefore, hyperparameter details are not relevant.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A] .
  34. Justification: This is a theoretical work; therefore, experimental statistical testing is not relevant.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [N/A] .
  39. Justification: This is a theoretical work; thus, information about computational resources is not applicable.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes] .
  44. Justification: We have reviewed the NeurIPS Code of Ethics Guidelines and verified that our paper fully complies with them.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A] .
  49. Justification: While explainability can have social implications, our work is theoretical in nature and therefore does not entail any direct social consequences.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A] .
  54. Justification: Our work is theoretical and therefore does not necessitate the implementation of safeguards.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [N/A] .
  59. Justification: This is a theoretical work; therefore, such assets are not relevant.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A] .
  64. Justification: This is a theoretical work; therefore, such assets are not relevant.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A] .
  69. Justification: This is a theoretical work; therefore, no crowdsourcing experiments were relevant.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A] .
  74. Justification: This is a theoretical study; therefore, no user studies were conducted.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A] .
  79. Justification: This is a theoretical work and does not involve any non-standard use of LLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix
The appendix provides formal definitions and proofs referenced throughout the paper.
  • Appendix A discusses extended related work.
  • Appendix B formalizes the GAMs that were studied in this work.
  • Appendix C provides additional details regarding the explainability queries analyzed in this work.
  • Appendix D contains possible extensions of our framework and its relationship to previous results.
  • Appendix E contains the proof of Proposition 1.
  • Appendix F contains the proof of Proposition 2.
  • Appendix G contains the proof of Proposition 3.
  • Appendix H contains the proof of Proposition 4.
  • Appendix I contains the proof of Proposition 5.
  • Appendix J contains the proof of Proposition 6.
  • Appendix K contains the proof of Proposition 7.
Appendix A Extended Related Work
In this section, we present a more technical overview of related work relevant to our framework. While all explanation queries we examine have been explored in prior studies, they were analyzed in the context of different models [15, 17, 2, 53, 54, 42, 102, 81, 49, 5, 65, 34, 82]. Notably, some results for neural networks [15, 2, 23] and tree ensembles [90, 54, 14, 21] are somewhat related to our setting, as NAMs and EBMs are instances of GAMs that incorporate neural networks or tree ensembles as base models. However, as demonstrated in our work (e.g., Table 2), the shift from additive to non-additive structures leads to fundamentally different complexity behaviors, necessitating entirely distinct proof techniques for both membership and hardness results.
Another line of related work involves studies on the computational complexity of generating specific types of explanations for linear models (e.g., [15, 99, 78, 102]). Some of the constructions used in our proofs draw upon ideas introduced in these works, and we reference them where appropriate — for instance, the use of sorting algorithms in the MSR/MCR queries and dynamic programming techniques for the CC and SHAP queries. That said, our study is the first to explore these explanations within the context of GAMs, while also addressing a substantially broader and more diverse set of input settings, component models, and explanation types — requiring us to tackle considerably more complex instances. For instance, proving hardness for various explanation queries over general discrete and continuous domains in models like NAMs and EBMs required novel constructions that differ substantially from those applicable to linear models.
Appendix B Generalized Additive Models
In this appendix, we outline the different generalized additive models (GAMs) analyzed in this work. Our focus will be on GAMs used for either regression or classification tasks. For simplicity, we assume the GAM produces a single output: a value in $\mathbb{R}$ for regression tasks or a single class from $\{0,1\}$ for classification tasks. However, our results can be extended to handle multiple regression outputs or multi-label classification. Specifically, we formalize a GAM for a regression task as follows:
$f(\textbf{x}):=\beta_{0}+\beta_{1}f_{1}(\textbf{x}_{1})+\beta_{2}f_{2}(\textbf{x}_{2})+\ldots+\beta_{k}f_{k}(\textbf{x}_{k})$ (3)
where $f_{1},\ldots,f_{k}$ represent uncorrelated functions $f_{i}:\mathcal{X}_{i}\to\mathbb{R}$, and $\beta_{0},\beta_{1},\beta_{2},\ldots,\beta_{k}\in\mathbb{Q}$ denote the intercept terms obtained during a training process. For classification tasks, we define similarly:
$f(\textbf{x}):=\text{step}\big(\beta_{0}+\beta_{1}f_{1}(\textbf{x}_{1})+\beta_{2}f_{2}(\textbf{x}_{2})+\ldots+\beta_{k}f_{k}(\textbf{x}_{k})\big)$ (4)
where we define the step function as $\text{step}(\textbf{z})=1$ if $\textbf{z}\geq 0$, and $\text{step}(\textbf{z})=0$ otherwise.
B.1 Component-Model Types
In this subsection, we define the component-model types discussed throughout the paper: (i) spline-based components within Smooth GAMs, (ii) neural network components within neural additive models (NAMs), and (iii) boosted tree components within explainable boosting machines (EBMs) .
Spline components for Smooth GAMs. Here, we define each $f_{i}$ as the standard piecewise spline function, corresponding to each component of the GAM [51]. Specifically, the domain $\mathcal{X}_{i}$ for each feature $\textbf{x}_{i}$ is divided into intervals $a_{1},a_{2}),[a_{2},a_{3}),\ldots,[a_{d-1},a_{d}]$, and a polynomial is defined for each interval $\mathcal{W}$. For $w\in\mathcal{W}$, we write the polynomial as $g(w):=\alpha_{r}w^{r}+\alpha_{r-1}w^{r-1}+\ldots+\alpha_{1}w+\alpha_{0}$. The function $f_{i}(\textbf{x}_{i})$ is defined over the domain $\mathcal{X}_{i}$ by assigning the value of the corresponding polynomial $g$ for each interval. Typically, the polynomials used in GAMs are of degree at most $3$ (referred to as cubic splines), which is the assumption we adopt here. We note, however, that many of our results also extend to higher-order polynomial splines. In particular, all of our hardness results remain valid in these cases, as such splines are at least as complex as cubic splines. As for the membership results, our findings for enumerable discrete settings still apply, since they rely only on the assumption that inference over each component can be performed in polynomial time — a condition satisfied by splines of any degree.
We also observe that in the simple case where each $f_{i}(\textbf{x}_{i})$ is the identity function — which is trivially captured by splines as well as our other models — the overall GAM $f(\textbf{x})$ effectively reduces to a standard linear model. This linear model can either be a regressor $f(\textbf{x}):=\beta_{0}+\beta_{1}\textbf{x}_{1}+\ldots+\beta_{k}\textbf{x}_{k}$ or a classification model $f(\textbf{x}):=\text{step}(\beta_{0}+\beta_{1}\textbf{x}_{1}+\ldots+\beta_{k}\textbf{x}_{k})$.
Neural network components for Neural Additive Models (NAMs). We denote a neural network $f$, consisting of t1t-1 hidden layers ($g^{j}$ where $j$ ranges from $1$ to t1t-1) and a single output layer ($g^{t}$). The layers are defined recursively — each layer $g^{(j)}$ is computed by applying the activation function $\sigma^{(j)}$ to the linear combination of the outputs from the previous layer $g^{(j-1)}$, the corresponding weight matrix $W^{(j)}$, and the bias vector $b^{(j)}$. This is represented as $g^{(j)}:=\sigma^{(j)}(g^{(j-1)}W^{(j)}+b^{(j)})$ for each $j$ in $1,\ldots,t$}. The model includes $t$ weight matrices ($W^{(1)},\ldots,W^{(t)}$), $t$ bias vectors ($b^{(1)},\ldots,b^{(t)}$), and $t$ activation functions ($\sigma^{(1)},\ldots,\sigma^{(t)}$).
[…]
Boosted tree ensemble components for Explainable Boosting Machines (EBMs) In this scenario, we assume that each individual model $f_{i}(\textbf{x}_{i})$ is a boosted tree ensemble. We begin by formalizing decision trees, followed by an extension to tree ensembles. A decision tree is an acyclic-directed graph that serves as a graphical model for a function $t:\mathcal{X}\to\mathbb{R}$ in regression and $t:\mathcal{X}\to\{0,1\}$ in classification. This graph represents the given function as follows: (i) Each internal node $v$ is associated with a unique decision rule, assumed to be of the form $\textbf{x}_{i}\geq r$ or $\textbf{x}_{i}<r$ for any $r\in\mathbb{Q}$, (ii) Every internal node $v$ has exactly two outgoing edges corresponding to the values $\{0,1\}$ assigned to $v$, where $1$ indicates the decision rule is satisfied, and 0 indicates it is not; (iii) Each leaf node is associated with a numerical value in $\mathbb{R}$ (for regression tasks) or with an associated class $j\in[c]$ (for classification tasks). Consequently, assigning a value to the input uniquely determines a path from the root to a leaf in the DT.
Regarding tree ensembles, each $f_{i}$ consists of an ensemble of $k^{i}\in\mathbb{N}$ decision trees (defined earlier) $t_{1},\ldots,t_{k^{i}}$, with each tree assigned a weight $\phi_{i}\in\mathbb{Q}$. In regression tasks, $f_{i}$ is expressed as a weighted sum of the individual tree outputs: $f_{i}(\textbf{x}_{i}):=\phi_{1}t_{1}(\textbf{x}_{i})+\phi_{2}t_{2}(\textbf{x}_{i})+\ldots+\phi_{k^{i}}t_{k^{i}}(\textbf{x}_{i})+\phi_{0}$. For classification, the prediction corresponds to the class receiving the highest weighted vote: $f_{i}(\textbf{x}_{i}):=\arg\max_{j\in[c]}\sum_{i=1}^{k^{i}}\phi_{i}\cdot\mathbb{I}[t(\textbf{x}_{i})=j]$, where $\mathbb{I}$ is the indicator function.
Appendix C Additional Query Formalizations
Here, we define the two remaining queries discussed throughout the paper: the Check-Sufficient-Reason (CSR) and Count Completions (CC) queries, both previously studied in [15, 17]. We then elaborate on the computational complexity of computing Shapley values, as examined in this work.
The CSR query determines whether a given subset $S$ qualifies as a sufficient reason. More formally:
Check Sufficient Reason (CSR): Input: A model $f$, an instance x, and a subset $S$. Output: Yes, if $S$ is a sufficient reason of $\langle f,\textbf{x}\rangle$, and No otherwise.
For the CC query, we consider a relaxed variant of the CSR query. Rather than checking whether a specific subset is sufficient, it measures the relative proportion of feature completions that preserve the original prediction, assuming the remaining features are independently and uniformly distributed. We begin by defining the completion count for a given subset:
$c(S,f,\textbf{x}):=\frac{{|\{\textbf{z}\in\mathcal{X}_{\bar{S}}\ |\ }{f(\textbf{x}_{S};\textbf{z}_{\bar{S}})=f(\textbf{x})\}|}}{{|\{\textbf{z}\in\{0,1\}^{|\overline{S}|}|}}$ (5)
Where $\mathcal{X}_{\bar{S}}$ denote the joint domain of the input features $\mathcal{X}_{i}$ for all $i\in\overline{S}$. The CC query is then defined as follows:
CC (Count Completions): Input: Model $f$, input x, and subset of features $S$. Output: The completion count $c(S,f,\textbf{x})$.
We present here a more detailed formalization of the Shapley value attribution method used in the paper. The Shapley value is defined as follows:
$\phi_{i}(f,\textbf{x}):=\sum_{S\subseteq[n]\setminus\{i\}}\frac{|S|!(n-|S|-1)!}{n!}(v(S\cup\{i\})-v(S))$ (6)
where $v(S)$ is the value function, and we use the common conditional expectation value function $v(S):=\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f(\textbf{z})|\textbf{z}_{S}=\textbf{x}_{S}]$ [100, 76]. We follow common conventions in frameworks that assessed the computational complexity of computing Shapley values [9, 102], as well as practical frameworks that compute Shapley values, such as the kernelSHAP method in the SHAP libary [76], and assume that each input feature is independent of all other features. In the discrete settings, every feature $i\in[n]$ is assigned some probability value $[0,1]$. These are called product distributions in the work of [9] or fully-factorized in the work of [102]. Formally, given some set of discrete values $[k]$, we can describe a probability function $p:[n]\times[k]\to[0,1]$. For example, $p(2,7)=\frac{1}{9}$, implies that the probability of feature $i=2$, to be set to the value $k=7$ is $\frac{1}{9}$. Then we can define $\mathcal{D}_{p}$ as an independent distribution over $\mathcal{X}$ iff:
$\textbf{Pr}(\textbf{x}):=\Big({\displaystyle\prod_{i\in[n],\textbf{x}_{i}=j}p(i,j)}\Big)$ (7)
The uniform distribution is a special case of $\mathcal{D}_{p}$, obtained by setting $p(i):=\frac{1}{k}$ for every $i\in[n]$. In the continuous setting, we define a probability function $p(\textbf{x}_{i})\in[0,1]$ for each feature $\textbf{x}_{i}$. Under the feature independence assumption, the joint probability for the continuous input setting is given by $\textbf{Pr}(\textbf{x}):=\prod_{i\in[n]}p(\textbf{x}_{i})$. Additionally, we assume that $p(\textbf{x}_{i})$ satisfies certain basic computational properties — specifically, that the sum $\sum_{j\in\mathcal{X}_{i}}p(i,j)$ can be computed in polynomial time (which we term poly-summability), or in the continuous case, that integration over $\mathcal{X}_{i}$ can be done in polynomial time (termed poly-enumerability). These conditions are satisfied by many common distributions, including the uniform distribution, where each input is assigned a fixed probability.
Appendix D Framework Extensions and Relationship to Previous Results
Regression, Probabilistic and Multi-Label Classification. In this segment, we discuss extensions of our framework to broader output configurations, relevant relaxations, and how these settings relate to existing works.
Our definitions can potentially be relaxed to incorporate probabilistic notions of sufficiency [104, 60, 6, 105], multi-output classification, or applications within bounded $\epsilon$-ball regions [109, 62, 77]. In particular, the framework could extend beyond binary classification to cover regression or probabilistic classification. For example, in a regression setting where $f:\mathbb{F}\to\mathbb{R}$, a sufficient reason may be defined as a subset $S\subseteq\{1,\ldots,n\}$ of input features such that:
[…]
for a given $\ell_{p}$-norm and some $\delta\in[0,1]$. Other notions discussed in our work — such as contrastive explanations and the definition of feature redundancy — can similarly be adapted to this framework. From a computational complexity standpoint, transitioning to regression for these queries does not change the underlying complexity, as the problem effectively reduces to a binary decision: whether the output is above or below some threshold $\delta$. Provided that $\delta$ is not fixed, hardness results naturally extend to cover the full domain. Likewise, membership results continue to rely on guessing witness assignments within this domain. This stands in sharp contrast to the unique SHAP query, whose behavior — and thus complexity — differs significantly between regression and classification settings, as shown in our work. The same reasoning extends to multi-output classification. In terms of hardness, results established for the single-output setting carry over directly. For membership results — whether they establish polynomial-time tractability or membership in higher complexity classes — it is important to note that they do not rely on the binary nature of inputs. Instead, they are based on guessing witness assignments and verifying conditions such as $f(\textbf{x}_{S};\textbf{z}_{\bar{S}})=f(\textbf{x})$ or $f(\textbf{x}_{S};\textbf{z}_{\bar{S}})\neq f(\textbf{x})$, which apply similarly in both single- and multi-output configurations.
The input setting impact on other ML models. In the main paper, we noted that while the input setting plays a significant role in shaping the computational complexity of explanation tasks in GAMs, prior studies on other ML models — specifically decision trees, neural networks, and tree ensembles — have not demonstrated a similar complexity separation based on the input domain. In this section, we provide a detailed justification for this claim, explaining why existing results suggest that such separations do not occur for these models. We begin with the simpler case of decision trees and tree ensembles. Intuitively, transitioning from a discrete to a continuous input domain does not alter the fundamental behavior of these models, as they compute piecewise constant functions (or an ensemble of such functions) and their split rules do not depend on specific input configurations. More concretely, prior work has shown that the complexity results for sufficient and contrastive explanations extend to continuous domains (see, e.g.,[54]), and similar extensions hold for counting-based explanations such as SHAP (e.g.,[55]). Lastly, since tree ensembles are simply aggregations of $k$ decision trees, the same proofs naturally extend to them as well.
In the context of neural networks, hardness results are already established even for Boolean input settings — this holds for the CSR, MCR, MSR, and CC queries [15], the FR query [17], and the SHAP query [102]. These results directly imply hardness in the enumerable discrete setting and, by extension, in the general discrete setting as well. Furthermore, [96] (building on [69]) demonstrated that satisfiability instances over binary inputs can be reduced to continuous input cases by introducing additional activation gadgets. This reduction, in turn, implies that the same hardness results hold for all these queries in the continuous setting.
Regarding membership results, prior work has already established that these results hold for binary inputs [15, 17, 102]. The same membership arguments — typically based on guessing a witness assignment — extend naturally to enumerable and general discrete domains, where the guess involves a $q$-bit vector instead of a binary input. Additionally, [96] demonstrate that verifying satisfiability queries for MLPs with ReLU activations over continuous inputs is in NP. The key idea is to guess the activation pattern of each ReLU and then solve a corresponding linear program, which is solvable in polynomial time. The CSR query (with $S=\emptyset$) effectively negates a satisfiability query, implying that CSR remains coNP-complete for continuous MLPs. Recall that the MSR query for MLPs is known to be $\Sigma^{P}_{2}$-complete [15], due to its use of a coNP oracle to check whether a subset is sufficient (i.e., solving the CSR query). Since CSR extends to continuous domains, the same reasoning carries over to MSR. Similar logic applies to the CC and SHAP queries, which are counting analogues of these reductions [15, 102, 8]. Altogether, these results underscore that for neural networks, the complexity class of each query remains stable across input domains, unlike the distinct separations seen in GAMs.
Appendix E Proof of Proposition 1
Proposition 1.
Given any GAM over an enumerable discrete domain, or a Smooth GAM over a discrete or continuous domain, the CSR, MCR, and MSR queries are solvable in polynomial time.
Proof. To start, we present the polynomial-time algorithms applicable to the enumerable discrete setting, which hold for any GAM. We then extend our discussion to Smooth GAMs over general discrete and continuous domains.
We divide the proof for the three explainability queries — CSR, MCR, and MSR — for GAMs over enumerable discrete settings into three separate lemmas, each establishing the corresponding complexity result. We note that the proofs across all these settings follow a similar structure to those used for linear models [15, 99, 78], with the main distinction being the need to account for the dependency on computing the minimum and maximum values of each component model $f_{i}(\textbf{x}_{i})$ to resolve the query.
Lemma 1.
The CSR query can be solved for a GAM $f$ over an enumerable discrete input space in polynomial time.
Proof. Let there be some instance $\langle f,\textbf{x},S\rangle$ where $f$ is a GAM. We note that given some feature $i$ we can evaluate its maximum negative affect (which we will call, similarly to [15], its ”penalty” by fixing it to the (negative) affecting edge of the domain $\mathcal{X}_{i}$. Formally, we define:
$p_{i}:=\begin{cases}\text{min}\{\beta_{i}\cdot f_{i}(\textbf{x}_{i})\ |\ \textbf{x}_{i}\in\mathcal{X}_{i}\}\quad if\ \ f(\textbf{x})>0\\ \text{max}\{\beta_{i}\cdot f(\textbf{x}_{i})\ |\ \textbf{x}_{i}\in\mathcal{X}_{i}\}\quad 0\quad otherwise\end{cases}$ (9)
Since the values in $\mathcal{X}_{i}$ are enumerable discrete, both $\text{min}\{\beta_{i}\cdot f_{i}(\textbf{x}_{i})\ |\ \textbf{x}_{i}\in\mathcal{X}_{i}\}$ as well as $\text{max}\{\beta_{i}\cdot f_{i}(\textbf{x}_{i})\ |\ \textbf{x}_{i}\in\mathcal{X}_{i}\}$ can be determined in polynomial time by simply iterating over any possible value of $\textbf{x}_{i}$. We can then compute all possible $p_{i}$ values in polynomial time. We can then simply check whether:
$\text{sign}(\sum_{i\in S}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\sum_{j\in\bar{S}}p_{j}+\beta_{0}\big)\cdot\text{sign}(\sum_{i\in n}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\beta_{0})\geq 0$ (10)
We recall that $(\sum_{i\in n}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\beta_{0})$ represents the value of the input x when passed through $f$ (prior to the step function), and $(\sum_{i\in S}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\sum_{j\in\bar{S}}p_{j}+\beta_{0})$ denotes the value for the input vector $(\textbf{x}_{S};\textbf{z}_{\bar{S}})$, where this value is maximally distant from the value of x when passed through $f$ for all $\textbf{z}\in\mathbb{F}$ (again, prior to the step function). Consequently, the signs of both sums align if and only if both are either positive or negative. This means the sign of the prediction $f(\textbf{x})$ before the step function matches the sign of $f(\textbf{x}_{S};\textbf{z}_{\bar{S}})$ before the step function. Since $(\textbf{x}_{S};\textbf{z}_{\bar{S}})$ represents the vector whose value through $f$ achieves the maximal distance from x when passed through $f$ (before the step function), the following equality holds if and only if, for any possible $\textbf{z}^{\prime}\in\mathbb{F}$, the predictions of $f$ over x and $(\textbf{x}_{S};\textbf{z}^{\prime}_{\bar{S}})$ are both positive (or both negative) before the step function. Therefore, after applying the step function, it holds that $\forall\textbf{z}^{\prime}\in\mathbb{F}$, $f(\textbf{x})=f(\textbf{x}_{S};\textbf{z}^{\prime}_{\bar{S}})$, which holds by definition if and only if $S$ is a sufficient reason concerning $\langle f,\textbf{x}\rangle$.
Lemma 2.
The MCR query can be solved for a GAM $f$ over an enumerable discrete input space in polynomial time.
Proof. Given an instance $\langle f,\textbf{x},k\rangle$, we use the same definition as in Lemma 1 of $p_{i}$ to denote the “penalty” of each feature $i$ in the GAM. We perform the following algorithm (Alg. 1), which computes a cardinally minimal contrastive reason. We assume, without loss of generality, that $f(\textbf{x})=1$. The alternate case, where $f(\textbf{x})=0$, can be addressed similarly by reversing Line 4, which will accordingly reverse the steps in the proof.
Input $f$, x
Algorithm 1 Cardinally Minimal Contrastive Reason Search
1:Compute all values $p_{i}$ for every $i\in[n]$
[…]
4:Sort $F$ in descending order by the value of $v_{i}:=\beta_{i}\cdot\textbf{x}_{i}-p_{i}$ for each $i$
5:for each $i\in F$ do
6:  if $\text{sign}(\sum_{i\in\bar{S}}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\sum_{j\in S}p_{j}+\beta_{0}\big)\cdot\text{sign}(\sum_{i\in n}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\beta_{0})<0$ then
7:   Break
8:  end if
9:  $S\leftarrow S\cup\{i\}$
10:end for
11:return $S$ $\triangleright$ $S$ is a cardinally minimal contrastive reason
We note that the computations of all $p_{i}$ values (Line 5 in Algorithm 1) can be computed in polynomial time due to our assumption of an enumerable discrete input space. The discrete input space means that for any $\textbf{x}_{i}$ we can simply iterate over the entire domain $\mathcal{X}_{i}$ in polynomial time and hence obtain both $\text{min}\{\beta_{i}\cdot f_{i}(\textbf{x}_{i})\ |\ \textbf{x}_{i}\in\mathcal{X}_{i}\}$ as well as $\text{max}\{\beta_{i}\cdot f_{i}(\textbf{x}_{i})\ |\ \textbf{x}_{i}\in\mathcal{X}_{i}\}$. Hence, it is clear that the algorithm runs in polynomial time.
To establish the correctness of our Lemma, it remains to demonstrate the validity of the algorithm:
Lemma 3.
Algorithm 1 obtains a cardinally minimal contrastive reason for $\langle f,\textbf{x}\rangle$.
Proof. We will first prove that algorithm 1 produces a valid contrastive reason for $\langle f,\textbf{x}\rangle$. In Lemma 1, we established that if the inequality $\text{sign}\left(\sum_{i\in S}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\sum_{j\in\bar{S}}p_{j}+\beta_{0}\right)\cdot\text{sign}\left(\sum_{i\in n}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\beta_{0}\right)\geq 0$ holds, then $S$ serves as a sufficient reason for $\langle f,\textbf{x}\rangle$. Similarly, if the relation $\text{sign}\left(\sum_{i\in\bar{S}}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\sum_{j\in S}p_{j}+\beta_{0}\right)\cdot\text{sign}\left(\sum_{i\in n}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\beta_{0}\right)\geq 0$ holds, then $\bar{S}$ is a sufficient reason, which implies by definition that $S$ is not a contrastive reason. Therefore, the following equivalence holds: if and only if $\text{sign}\left(\sum_{i\in S}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\sum_{j\in\bar{S}}p_{j}+\beta_{0}\right)\cdot\text{sign}\left(\sum_{i\in n}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\beta_{0}\right)<0$, then $S$ is a contrastive reason with respect to $\langle f,\textbf{x}\rangle$.
We will now demonstrate that the generated set $S$ is a cardinally minimal contrastive reason with respect to $\langle f,\textbf{x}\rangle$. Let $1\leq\ell\leq n$ represent the last feature added to $S$ in line 10 of algorithm 1. Then, for $S^{\prime}:=S\setminus\{\ell\}$, it follows that: $\text{sign}\left(\sum_{i\in S^{\prime}}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\sum_{j\in\bar{S^{\prime}}}p_{j}+\beta_{0}\right)\cdot\text{sign}\left(\sum_{i\in n}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\beta_{0}\right)\geq 0$, implying that $S^{\prime}$ is not a contrastive reason for $\langle f,\textbf{x}\rangle$.
Recall that $F$ was sorted in descending order according to the values $v_{i}:=\beta_{i}\cdot f_{i}(\textbf{x}_{i})-p_{i}$, and $S^{\prime}$ therefore represents the subset of size $|S^{\prime}|$ that maximizes $\sum_{j\in S^{\prime}}(\beta_{j}\cdot f_{j}(\textbf{x}_{j})-p_{j})$ and hence also maximizes the difference
$(\sum_{i\in n}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\beta_{0})-(\sum_{i\in\bar{S^{\prime}}}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\sum_{j\in S^{\prime}}p_{j}+\beta_{0})=\sum_{j\in\bar{S^{\prime}}}(\beta_{j}\cdot f_{j}(\textbf{x}_{j})-p_{j})$ (11)
This implies that any subset $S^{\prime\prime}\subseteq[n]$ of size $|S^{\prime\prime}|\geq|S^{\prime}|$ satisfies: $\text{sign}\left(\sum_{i\in S^{\prime\prime}}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\sum_{j\in\bar{S^{\prime\prime}}}p_{j}+\beta_{0}\right)\cdot\text{sign}\left(\sum_{i\in n}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\beta_{0}\right)\geq 0$, indicating that $S^{\prime\prime}$ is not a contrastive reason for $\langle f,\textbf{x}\rangle$. Overall, we have shown that $S$ is a contrastive reason for $\langle f,\textbf{x}\rangle$, and that any subset $S^{\prime\prime}$ of size $|S^{\prime\prime}|\geq|S^{\prime}|$ is not a contrastive reason for $\langle f,\textbf{x}\rangle$. By definition, $|S^{\prime}|=|S|-1$, so we have established that there is no subset of size |S|1|S|-1 or smaller that is a contrastive reason for $\langle f,\textbf{x}\rangle$. This confirms that $S$ is a cardinally minimal contrastive reason for $\langle f,\textbf{x}\rangle$.
Having established that algorithm 1 produces a cardinally minimal contrastive reason for $S$, solving the MCR query in polynomial time simply requires running the algorithm and verifying if $|S|\geq k$, thereby completing our proof.
Lemma 4.
The MSR query can be solved for a GAM $f$ over an enumerable discrete input space in polynomial time.
Proof. Given an instance $\langle f,\textbf{x},k\rangle$, we will show that the MSR query can be resolved in polynomial time. For a specific feature $i$, we will again (as in Lemma 1) refer to its "penalty" as the maximum negative impact it can exert on the prediction $f(\textbf{x})$. Formally, we define $p_{i}:=\text{min}\{\beta_{i}\cdot f_{i}(\textbf{x}_{i})\ |\ \textbf{x}_{i}\in\mathcal{X}_{i}\}$ if $f(\textbf{x})>0$, and $p_{i}:=\text{max}\{\beta_{i}\cdot f_{i}(\textbf{x}_{i})\ |\ \beta_{i}\in\mathcal{X}_{i}\}$ otherwise. Furthermore, we require a measure for the value of feature $i$ when it is fixed. We define $v_{i}:=\beta_{i}\cdot f_{i}(\textbf{x}_{i})-p_{i}$. Note that when $\sum_{i\in[n]}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\beta_{0}<0$, i.e., when $f(\textbf{x})=0$, a “good score” will be a highly negative one. Thus, we specifically define $v_{i}:=(\beta_{i}\cdot f_{i}(\textbf{x}_{i})-p_{i})\cdot\text{sign}(\sum_{i\in[n]}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\beta_{0})$. However, for simplicity, we will assume without loss of generality that $f(\textbf{x})>0$, and therefore $v_{i}:=\beta_{i}\cdot f_{i}(\textbf{x}_{i})-p_{i}$. Intuitively, $v_{i}$ reflects the value of feature $i$ when it is held constant rather than allowed to take on any value, thereby influencing the penalty score $p_{i}$. With these definitions, we are prepared to propose the following algorithm for obtaining a cardinally minimal sufficient reason for the given instance $\langle f,\textbf{x}\rangle$.
Input $f$, x
Algorithm 2 Cardinally Minimal Sufficient Reason Search
1:Compute all values $v_{i},p_{i}$ for every $i\in[n]$
[…]
4:Sort $F$ in descending order by the value of $v_{i}:=\beta_{i}\cdot f_{i}(\textbf{x}_{i})-p_{i}$ for each $i$
5:for each $i\in F$ do
6:  if $\text{sign}(\sum_{i\in S}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\sum_{j\in\bar{S}}p_{j}+\beta_{0}\big)\cdot\text{sign}(\sum_{i\in n}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\beta_{0})\geq 0$ then
7:   Break
8:  end if
9:  $S\leftarrow S\cup\{i\}$
10:end for
11:return $S$ $\triangleright$ $S$ is a cardinally minimal sufficient reason
Lemma 5.
Algorithm 2 obtains a cardinally minimal sufficient reason for $\langle f,\textbf{x}\rangle$.
Proof. First let us prove that Algorithm 2 provides a valid sufficient reason. This result is straightforward since we have proven in Lemma 1 that $S$ is a sufficient reason of $\langle f,\textbf{x}\rangle$ iff the condition $\text{sign}(\sum_{i\in\bar{S}}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\sum_{j\in S}p_{j}+\beta_{0}\big)\cdot\text{sign}(\sum_{i\in n}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\beta_{0})\geq 0$ holds.
We will now demonstrate that the generated set $S$ is a cardinally minimal sufficient reason with respect to $\langle f,\textbf{x}\rangle$. Let $1\leq\ell\leq n$ represent the last feature added to $S$ in line 10 of algorithm 2. Then, for $S^{\prime}:=S\setminus\{\ell\}$, it follows that: $\text{sign}\left(\sum_{i\in S^{\prime}}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\sum_{j\in\bar{S^{\prime}}}p_{j}+\beta_{0}\right)\cdot\text{sign}\left(\sum_{i\in n}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\beta_{0}\right)<0$, implying that $S^{\prime}$ is not a sufficient reason for $\langle f,\textbf{x}\rangle$.
Recall that $F$ was sorted according to the values $v_{i}:=\beta_{i}\cdot f_{i}(\textbf{x}_{i})-p_{i}$ in descending order, and $S^{\prime}$ therefore represents the subset of size $|S^{\prime}|$ that maximizes $\sum_{j\in S^{\prime}}(\beta_{j}\cdot f_{j}(\textbf{x}_{j})-p_{j})$ and hence $\overline{S^{\prime}}$ represent the subsets that minimizes the difference $\sum_{j\in\bar{S^{\prime}}}(\beta_{j}\cdot f_{j}(\textbf{x}_{j})-p_{j})$. This is equivalent to stating that $\overline{S^{\prime}}$ represents the set of size $|\overline{S^{\prime}}|$ that minimizes:
$(\sum_{i\in S^{\prime}}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\sum_{j\in\bar{S^{\prime}}}p_{j}+\beta_{0})-(\sum_{i\in n}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\beta_{0})=\sum_{j\in\bar{S^{\prime}}}(\beta_{j}\cdot f_{j}(\textbf{x}_{j})-p_{j})$ (12)
This implies that any subset $S^{\prime\prime}\subseteq[n]$ for which $|S^{\prime\prime}|<|S^{\prime}|$ and hence also $|\overline{S^{\prime\prime}}|\geq|\overline{S^{\prime}}|$ satisfies: $\text{sign}\left(\sum_{i\in S^{\prime\prime}}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\sum_{j\in\bar{S^{\prime\prime}}}p_{j}+\beta_{0}\right)\cdot\text{sign}\left(\sum_{i\in n}\beta_{i}\cdot f_{i}(\textbf{x}_{i})+\beta_{0}\right)<0$, indicating that $\overline{S^{\prime\prime}}$ is a contrastive reason for $\langle f,\textbf{x}\rangle$, and hence that $S^{\prime\prime}$ is not a sufficient reason for $\langle f,\textbf{x}\rangle$. Overall, we have shown that $S$ is a sufficient reason for $\langle f,\textbf{x}\rangle$, and that any subset $S^{\prime\prime}$ of size $|S^{\prime\prime}|<|S^{\prime}|$ is not a sufficient reason for $\langle f,\textbf{x}\rangle$. By definition, $|S^{\prime}|=|S|-1$, so we have established that there is no subset of size |S|1|S|-1 or smaller that is a sufficient reason for $\langle f,\textbf{x}\rangle$. This confirms that $S$ is a cardinally minimal sufficient reason for $\langle f,\textbf{x}\rangle$.
Having established that algorithm 2 produces a cardinally minimal sufficient reason for $S$, solving the MSR query in polynomial time simply requires running the algorithm and verifying if $|S|\geq k$, thereby completing our proof.
Having established the complexity results for the three explanation types in the enumerable discrete input setting, we now turn our attention to the general discrete or continuous input setting, focusing on the case of Smooth GAMs. In fact, we will show that this claim holds for a broader class of model families, including other tractable components like decision trees:
Lemma 6.
Given a Smooth GAM over a general discrete or continuous input setting —- then the CSR, MCR and MSR queries can be obtained in polynomial time.
Proof. We have established in Lemma 1, Lemma 2, and Lemma 4 that, assuming the minimum and maximum viable values for each model component $f_{i}$ can be determined within the domain $\mathcal{X}_{i}$, the CSR, MSR, and MCR queries can be computed in polynomial time. We will now demonstrate this for each corresponding component model.
First, consider the case where each $f_{i}$ is the identity function (i.e., $f$ is a linear model). In this case, determining the minimum and maximum values in each domain is straightforward and can be done in polynomial time. For cubic splines, this result extends naturally, as it involves iterating over each defined region of the piecewise polynomial $f_{i}$. For each region, we evaluate the value of $f_{i}$ at the domain edges and at the extreme points, which can be found by solving a polynomial equation of degree 3 (achievable in polynomial time). Among these values (edges and polynomial extremes), the minimum and maximum are selected. If $\mathcal{X}_{i}$ is a general discrete set, we iterate over the continuous region representing $\mathcal{X}_{i}$, find the extreme points as before, and for each extreme point, identify the two nearest discrete points within $\mathcal{X}_{i}$ — one above and one below. From this set of viable extreme points of the discrete domain $\mathcal{X}_{i}$, we then select the minimum and maximum values.
We note that this proof applies not only to Smooth GAMs but to any model where the minimum and maximum attainable values of each $f_{i}$ can be computed in polynomial time. This includes, for example, decision trees. In the case where $f_{i}$ is a decision tree, we iterate over all leaf edges (representing all viable values of $f_{i}$) and select the minimum and maximum values from these leaves. This process can also be completed in polynomial time with respect to the size of the tree.
Appendix F Proof of Proposition 2
Proposition 2.
Given a Smooth GAM $f$ over a continuous input space, then the FR query can be obtained in polynomial time.
Proof. We will prove this for a more general class that includes any GAM $f$ defined over continuous inputs, where the output $f(\mathbf{x})$ is continuous, and for which both $\text{min}\{\beta_{i}f(i)\}$ and $\text{max}\{\beta_{i}f(i)\}$ can be computed in polynomial time for any feature $i$. The polynomial-time computation of $\text{min}\{\beta_{i}f(i)\}$ and $\text{max}\{\beta_{i}f(i)\}$ for splines has been proven in Lemma 1. Furthermore, since splines are defined as piecewise polynomials, it is clear that $f(\mathbf{x})$ is a continuous function. Therefore, proving the complexity results for the FR query for this family of classifiers will also encompass GAMs defined over splines.
We first note that by definition a feature $i$ is redundant with respect to $f$ if and only if for any $\textbf{x}^{\prime},\textbf{x}_{i},\textbf{z}_{i}$ it holds that $f(\textbf{x}^{\prime}_{[n]\setminus\{i\}};\textbf{x}_{i})=f(\textbf{x}^{\prime}_{[n]\setminus\{i\}};\textbf{z}_{i})$ (or in other words, changing the value of feature $i$ cannot cause the classification to change. We will prove the following claim:
Lemma 7.
Let there be some GAM $f$ where both the input domain $\mathcal{X}$ is continuous, and the output $f(\textbf{x})$ is continuous. Let us assume that $i$ is redundant with respect to $\langle f\rangle$ and that $f$ is not trivial (in other words, a function that always outputs $1$ or always outputs 0). Then, it holds that for all $\textbf{x}_{i}\in\mathcal{X}_{i}$, $\beta_{i}f_{i}(\textbf{x}_{i})=0$.
Proof. First we recall that by definition $i$ is redundant with respect to $f$ iff for any $\textbf{x}^{\prime},\textbf{x}_{i},\textbf{z}_{i}$ it holds that $f(\textbf{x}^{\prime}_{[n]\setminus\{i\}};\textbf{x}_{i})=f(\textbf{x}^{\prime}_{[n]\setminus\{i\}};\textbf{x}_{i})$. In other words, this implies that:
$\wedge$ (13)
$`[\beta_{0}+\beta_{1}f_{1}(\textbf{x}_{1})+\ldots+f_{k}(\textbf{x}_{k})\geq 0]]$
$\vee[[\beta_{0}+\beta_{1}f_{1}(\textbf{x}_{1})+\ldots+\beta_{i-1}f_{i-1}(\textbf{x}_{i-1})+\beta_{i+1}f_{i+1}(\textbf{x}_{i+1})+\ldots+\beta_{k}f_{k}(\textbf{x}_{k})<0]\wedge$
$`[\beta_{0}+\beta_{1}f_{1}(\textbf{x}_{1})+\ldots+f_{k}(\textbf{x}_{k})<0]]$
Let us assume for contradiction that for any x, the term $\beta_{0}+\beta_{1}f_{1}(\textbf{x}_{1})+\ldots+\beta_{i-1}f_{i-1}(\textbf{x}_{i-1})+\beta_{i+1}f_{i+1}(\textbf{x}_{i+1})+\ldots+\beta_{k}f_{k}(\textbf{x}_{k})$ is always strictly positive or always strictly negative. Since $i$ is redundant, it follows that adding the term $\beta_{i}f_{i}(\textbf{x}_{i})$ to the rest of the sum cannot cause the classification to “flip,” meaning the overall term cannot change from positive to negative or vice versa. This implies that the entire sum $\beta_{0}+\beta_{1}f_{1}(\textbf{x}_{1})+\ldots+\beta_{k}f_{k}(\textbf{x}_{k})$ for any x is always positive or always negative, which contradicts the assumption that $f$ is non-trivial. Therefore, there must necessarily exist some $\textbf{x},\textbf{z}\in\mathcal{X}$ such that $\beta_{0}+\beta_{1}f_{1}(\textbf{x}_{1})+\ldots+\beta_{i-1}f_{i-1}(\textbf{x}_{i-1})+\beta_{i+1}f_{i+1}(\textbf{x}_{i+1})+\ldots+\beta_{k}f_{k}(\textbf{x}_{k})$ is positive, and $\beta_{0}+\beta_{1}f_{1}(\textbf{z}_{1})+\ldots+\beta_{i-1}f_{i-1}(\textbf{z}_{i-1})+\beta_{i+1}f_{i+1}(\textbf{z}_{i+1})+\ldots+\beta_{k}f_{k}(\textbf{z}_{k})$ is negative (or vice versa).
Moreover, since $\mathcal{X}$ is continuous, we can select x and z such that $\beta_{0}+\beta_{1}f_{1}(\textbf{x}_{1})+\ldots+\beta_{i-1}f_{i-1}(\textbf{x}_{i-1})+\beta_{i+1}f_{i+1}(\textbf{x}_{i+1})+\ldots+\beta_{k}f_{k}(\textbf{x}_{k})$ and $\beta_{0}+\beta_{1}f_{1}(\textbf{z}_{1})+\ldots+\beta_{i-1}f_{i-1}(\textbf{z}_{i-1})+\beta_{i+1}f_{i+1}(\textbf{z}_{i+1})+\ldots+\beta_{k}f_{k}(\textbf{z}_{k})$ are infinitesimally close to the decision boundary. We will denote these points as either $0^{+}$ or $0^{-}$, representing an infinitesimally close value that is either above or below 0. Now, assuming there exists some $\textbf{x}^{\prime}_{i}$ such that $\beta_{i}f_{i}(\textbf{x}^{\prime}_{i})$ is negative, we can take an infinitesimally small value for $\beta_{0}+\beta_{1}f_{1}(\textbf{x}_{1})+\ldots+\beta{i-1}f_{i-1}(\textbf{x}_{i-1})+\beta_{i+1}f_{i+1}(\textbf{x}_{i+1})+\ldots+\beta_{k}f_{k}(\textbf{x}_{k})$ above the decision boundary ($0^{+}$) such that adding $\beta_{i}f_{i}(\textbf{x}^{\prime}_{i})$ causes the decision to become negative. The analogous condition holds for the opposite direction of the decision boundary (assuming there exists some $\textbf{x}^{\prime}_{i}$ for which $\beta_{i}f_{i}(\textbf{x}^{\prime}_{i})$ is strictly negative). Consequently, the only way for $i$ to be redundant is if, for every $\textbf{x}^{\prime}_{i}\in\mathcal{X}_{i}$, it holds that $\beta_{i}f_{i}(\textbf{x}^{\prime}_{i})=0$, thereby concluding the proof of the lemma.
We will now use the previous Lemma to describe a simple polynomial-time algorithm for solving this task. The algorithm iterates over all features $i$ from $1$ to $n$. Since $f_{i}$ is a spline, we can compute the minimal and maximal values $\text{min}{\beta_{i}f_{i}(\textbf{x}_{i})}$ and $\text{max}{\beta_{i}f_{i}(\textbf{x}_{i})}$ in polynomial time. If either both terms are equal to 0 or $\beta_{i}=0$, we return that $i$ is redundant. Otherwise, we return that it is not. The correctness follows from the correctness of Lemma 7, the continuity of $\mathcal{X}$, and the continuity of $f(\textbf{x})$ (a result of working with splines). This concludes the proof.
Appendix G Proof of Proposition 2
Proposition 3.
Given any regression GAM with enumerable discrete inputs, or given a regression Smooth GAM with discrete or continuous inputs, then SHAP query is solvable in polynomial time.
Proof. We divide the proof into three separate lemmas. First, we establish the claim for general GAMs under the enumerable discrete setting. Next, we address Smooth GAMs over general discrete domains. Finally, we prove the claim for Smooth GAMs over continuous domains.
Lemma 8.
Assuming $f$ is a regression GAM over an enumerable discrete input space $\mathbb{F}$, then solving SHAP for $\langle f,\textbf{x},i\rangle$ can be solved in polynomial time.
Proof. For a certain value function $v(S)$, given the assumption of feature independence and the additive structure of $f$, we have:
$v(S)=\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f(\textbf{z})|\textbf{z}_{S}=\textbf{x}_{S}]=$ (14)
$\mathbb{E}_{\textbf{z}_{\bar{S}}}[f(\textbf{x}_{S};\textbf{z}_{\bar{S}})]=$
$\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f(\textbf{x}_{S};\textbf{z}_{\bar{S}})]=$
$\sum_{\textbf{z}\in\mathbb{F}}\Big({\displaystyle\prod_{\ell\in[n],\textbf{z}_{\ell}=j}p(\ell,j)}\Big)\cdot f(\textbf{x}_{S};\textbf{z}_{\bar{S}})=$
$\sum_{\textbf{z}\in\mathbb{F}}\Big({\displaystyle\prod_{\ell\in[n],\textbf{z}_{\ell}=j}p(\ell,j)}\Big)\cdot(\sum_{\ell\in S}\beta_{\ell}f_{\ell}(\textbf{x}_{\ell})+\sum_{j\in\bar{S}}\beta_{\ell}f_{j}(\textbf{z}_{j})+\beta_{0})=$
$\sum_{\ell\in S}\beta_{\ell}f_{\ell}(\textbf{x}_{\ell})+\sum_{j\in\bar{S}}\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[\beta_{j}f_{j}(\textbf{z}_{j})]+\beta_{0}$
Now, we have that:
$v(S\cup\{i\})-v(S)=$ (15)
$\big(\sum_{\ell\in S\cup\{l\}}\beta_{\ell}f_{\ell}(\textbf{x}_{\ell})+\sum_{j\in\bar{S}\setminus\{i\}}\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[\beta_{j}f_{j}(\textbf{z}_{j})]+\beta_{0}\big)$-
$\big(\sum_{\ell\in S}\beta_{\ell}f_{\ell}(\textbf{x}_{\ell})+\sum_{j\in\bar{S}}\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[\beta_{j}f_{j}(\textbf{z}_{j})]+\beta_{0}\big)=$
$\beta_{i}f_{i}(\textbf{x}_{i})-\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[\beta_{i}f_{i}(\textbf{x}_{i})]=$
$\beta_{i}(f_{i}(\textbf{x}_{i})-\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}(f_{i}(\textbf{z}_{i}))$
By incorporating it into the SHAP formulation, we obtain:
$\phi_{i}(f,\textbf{x})=\sum_{S\subseteq[n]\setminus\{i\}}\frac{|S|!(n-|S|-1)!}{n!}(v(S\cup\{i\})-v(S))=$ (16)
$\sum_{S\subseteq[n]\setminus\{i\}}\frac{|S|!(n-|S|-1)!}{n!}(\beta_{i}f_{i}(\textbf{x}_{i})-\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[\beta_{i}f_{i}(\textbf{x}_{i})])=$
$\frac{\beta_{i}}{n}(f_{i}(\textbf{x}_{i})-\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}(f_{i}(\textbf{z}_{i})))$
where the last equation follows from the fact that the inner product does not depend on $S$, and from the well-known result that the sum of the Shapley values, $\sum_{S\subseteq[n]\setminus\{i\}}\frac{|S|!(n-|S|-1)!}{n!}$, simplifies to $\frac{1}{n}$ through a telescoping sum. By definition, $f_{i}(\textbf{x}_{i})$ can be computed in polynomial time. Furthermore, since the domain $\mathcal{X}_{i}$ is enumerable discrete, $\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f_{i}(\textbf{z}_{i})]$ can also be computed in polynomial time by evaluating $\sum_{j\in\mathcal{X}_{i}}p(i,j)f_{i}(\textbf{z}_{i})$ (where the equivalence holds again by utilizing the feature independence property), thereby completing the proof.
Lemma 9.
Given a GAM $f$ over a general discrete input space, obtaining the SHAP query for regression can be solved in polynomial time.
Proof. We note that until our last step — proving tractability for the enumerable discrete setting in Lemma 8, which involved showing that $\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f_{i}(\textbf{z}_{i})]$ can be computed in polynomial time (by evaluating $\sum_{j\in\mathcal{X}_{i}}p(i,j)f_{i}(\textbf{z}_{i})$) — we did not assume explicit enumerability. Thus, all results up to this point also apply to the general discrete case. In the general setting, correctness follows directly from the poly-summability property of the distribution $\mathcal{D}_{p}$, thereby establishing tractability.
Lemma 10.
Given a GAM $f$ over a continuous input space with poly-integrable components, obtaining the SHAP for regression can be solved in polynomial time.
Proof. The proof will be similar to the discrete scenario. For a certain value function $v(S)$, given the assumption of feature independence and the additive structure of $f$, we have:
$v(S)=\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f(\textbf{z})|\textbf{z}_{S}=\textbf{x}_{S}]=$ (17)
$\mathbb{E}_{\textbf{z}_{\bar{S}}}[f(\textbf{x}_{S};\textbf{z}_{\bar{S}})]=$
$\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f(\textbf{x}_{S};\textbf{z}_{\bar{S}})]=$
$\int_{\textbf{z}\in\mathcal{X}}p(\textbf{z})\cdot f(\textbf{x}_{S};\textbf{z}_{\bar{S}})=$
$\int_{\textbf{z}\in\mathcal{X}}\big(p(\textbf{z})\cdot(\sum_{\ell\in S}\beta_{\ell}f_{\ell}(\textbf{x}_{\ell})+\sum_{j\in\bar{S}}\beta_{\ell}f_{j}(\textbf{z}_{j})+\beta_{0})\big)=$
$\sum_{\ell\in S}\beta_{\ell}f_{\ell}(\textbf{x}_{\ell})+\sum_{j\in\bar{S}}\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[\beta_{j}f_{j}(\textbf{z}_{j})]+\beta_{0}$
Now, the following condition holds:
$v(S\cup\{i\})-v(S)=$ (18)
$\big(\sum_{\ell\in S\cup\{l\}}\beta_{\ell}f_{\ell}(\textbf{x}_{\ell})+\sum_{j\in\bar{S}\setminus\{i\}}\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[\beta_{j}f_{j}(\textbf{z}_{j})]+\beta_{0}\big)$-
$\big(\sum_{\ell\in S}\beta_{\ell}f_{\ell}(\textbf{x}_{\ell})+\sum_{j\in\bar{S}}\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[\beta_{j}f_{j}(\textbf{z}_{j})]+\beta_{0}\big)=$
$\beta_{i}f_{i}(\textbf{x}_{i})-\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[\beta_{i}f_{i}(\textbf{x}_{i})]=$
$\beta_{i}(f_{i}(\textbf{x}_{i})-\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}(f_{i}(\textbf{z}_{i}))$
Incorporating it into the SHAP formulation yields:
$\phi_{i}(f,\textbf{x})=\sum_{S\subseteq[n]\setminus\{i\}}\frac{|S|!(n-|S|-1)!}{n!}(v(S\cup\{i\})-v(S))=$ (19)
$\sum_{S\subseteq[n]\setminus\{i\}}\frac{|S|!(n-|S|-1)!}{n!}(\beta_{i}f_{i}(\textbf{x}_{i})-\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[\beta_{i}f_{i}(\textbf{x}_{i})])=$
$\frac{\beta_{i}}{n}(f_{i}(\textbf{x}_{i})-\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}(f_{i}(\textbf{z}_{i})))$
The final equality follows from the fact that the inner product is independent of $S$, along with the well-known identity that the sum of Shapley coefficients, $\sum_{S\subseteq[n]\setminus{i}}\frac{|S|!(n-|S|-1)!}{n!}$, evaluates to $\frac{1}{n}$ via a telescoping argument. By definition, $f_{i}(\textbf{x}_{i})$ is computable in polynomial time. Thus, the question of whether the SHAP value is polynomial-time computable reduces to whether $\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f_{i}(\textbf{z}_{i})]$ can be computed in polynomial time. Under the feature independence assumption, we observe that:
$\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f_{i}(\textbf{z}_{i})]:=\int_{\textbf{z}_{i}\in\mathcal{X}_{i}}p(\textbf{z})f_{i}(\textbf{z}_{i})=\int_{\textbf{z}_{i}\in\mathcal{X}_{i}}p_{i}(\textbf{z}_{i})f_{i}(\textbf{z}_{i})$ (20)
Since $f_{i}(\textbf{z}_{i})$ is a smooth function bounded by degree 3, and $p(\textbf{z}_{i})$ is poly-integrable, it follows by definition that $\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f_{i}(\textbf{z}_{i})]$ can be computed in polynomial time by evaluating the integral $\int_{\textbf{z}_{i}\in\mathcal{X}_{i}}p_{i}(\textbf{z}_{i})f_{i}(\textbf{z}_{i})$. This completes the proof.
Appendix H Proof of Proposition 4
Proposition 4.
Given any unary-encoded GAM over a discrete, enumerable input domain with polynomially bounded output, the CC and SHAP-C queries admit a pseudo-polynomial-time algorithm.
We will divide the proof into two lemmas, addressing the CC and SHAP queries separately, beginning with the CC case.
Lemma 11.
Given any unary-encoded GAM over a discrete, enumerable input domain with polynomially bounded output, the CC query admits a pseudo-polynomial-time algorithm.
Proof. We prove this by reducing the CC problem when weights are provided in unary to a variant of the counting version of the classic #Knapsack problem, which we refer to as #Multi-Choice Knapsack. We then show that #Multi-Choice Knapsack can be solved via a dynamic programming algorithm. Our proof follows a similar structure to that of [15], who addressed linear models over binary domains, but introduces additional complexity to handle arbitrary enumerable discrete input configurations. In particular, whereas the proof in [15] addresses this problem via a reduction to #Knapsack, our approach reduces it to the #Multi-Choice Knapsack setting.
#\#Multi-Choice Knapsack: Input: $w:=(w_{1},\ldots,w_{n})$ where each $w_{i}$ is a non-negative integer weight, a vector of sets $(S_{1},\ldots,S_{n})$ where each set $S_{i}$ contains a fixed number of associated integers, and a target integer $C$ which denotes the capacity. Output: $|\{(x_{1},\ldots,x_{n})\in S_{1}\times\ldots\times S_{n}\ |\ \sum_{i\leq n}w_{i}x_{i}\leq C\}|$
We observe that this problem resembles the classic #\#Knapsack problem — the counting variant of Knapsack — except that each variable $x_{i}$ ranges over a set $S_{i}$ of enumerable integers, rather than being restricted to $\{0,1\}$. This range of possible values corresponds to the outputs of the GAM. Since we assume that the GAM uses an enumerably discrete unary input encoding and that its outputs are polynomially bounded, it follows that the set of possible outputs is itself polynomially bounded with respect to the input encoding (and therefore also unary). The work of [15] reduces the CC query for Perceptrons (i.e., linear models over binary inputs with unary weights) to the classic #\#Knapsack problem via a polynomial-time reduction, where $C$ is set to the bias $b$ of the linear model. We observe that applying the same reduction — except replacing binary inputs with sets of enumerable discrete values representing the enumerable set of unary-encoded possible outputs of the GAM — yields an instance of the #\#Multi-Choice Knapsack problem instead (and the target integer $C$ equals the intercept $\beta_{0}$ of the GAM instead of the bias $b$ of the linear model). Thus, it remains to show that #\#Multi-Choice Knapsack is solvable in pseudo-polynomial time, which we do by adapting the standard dynamic programming algorithm typically used for this problem (and also discussed in [15]).
The algorithm will be constructed as follows: given an instance of #Multi-Choice Knapsack which includes $w:=(w_{1},\ldots,w_{n})$, $(S_{1},\ldots,S_{n})$, and $C$, we can define the following quantity:
$DP[i][C]:=|\{(x_{1},\ldots,x_{n})\in S_{1}\times\ldots\times S_{n}\ |\ \sum_{i\leq n}w_{i}x_{i}\leq C\}|$ (21)
The final result for the #Multi-Choice Knapsack (and consequently for the CC query) is given by the value of $DP[n][b]$. This can be computed using an iterative dynamic programming approach based on the following inductive step:
$DP[i+1][C]=DP[i][C]+\sum_{sinS_{i+1}}DP[i][C-s]$ (22)
We carry out this iterative computation starting from the standard dynamic programming base case (as in the classic Knapsack and #Knapsack algorithms, and also noted in [15]): $DP[0][\alpha]=0$ for all $\alpha<0$ and $DP[0][\alpha]=1$ for all $\alpha\geq 0$. Given the pseudo-polynomial nature of the problem, a polynomial number of iterations suffices to compute the final value $DP[n][b]$, which yields the solution to the #Multi-Choice Knapsack and, consequently, the answer to the CC query in this setting.
Lemma 12.
Given any unary-encoded GAM over a discrete, enumerable input domain with polynomially bounded output, the SHAP query admits a pseudo-polynomial-time algorithm.
Proof. We build on the proof introduced by [102] which showed that the computation of SHAP for a model $f$ and some input x can be reduced to the computation of $\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f(\textbf{z})]$ in polynomial time, given that we assume feature independence. Hence, we are only left to prove that $\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f(\textbf{z})]$ can be obtained in pseudo-polynomial time for GAMs over enumerable discrete input spaces. We note that the following holds:
$\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[f(\textbf{z})]=\mathbb{E}_{\textbf{z}\sim\mathcal{D}_{p}}[\text{step}(\beta_{0}+\beta_{1}\cdot f_{1}(\textbf{z}_{1})+\ldots\beta_{n}\cdot f_{n}(\textbf{z}_{n}))]$ (23)
Since we assume that the input space $\mathcal{X}$ is enumerable discrete, computing $\mathbb{E}_{\textbf{z}\sim\mathcal{D}p}[f(\textbf{z})]$ reduces to counting the number of assignments $(x_{1},\ldots,x_{n})\in S_{1}\times\ldots\times S_{n}$ such that $\beta_{0}+\beta_{1}\cdot f_{1}(\textbf{z}_{1})+\ldots+\beta_{n}\cdot f_{n}(\textbf{z}_{n})\geq 0$. This is equivalent to counting assignments where $\sum{i=1}^{n}\beta_{i}\cdot f_{i}(\textbf{z}_{i})\geq-\beta_{0}$. Setting the threshold $C:=\beta_{0}$, the problem becomes an instance of the #Multi-Choice Knapsack problem described in Lemma 11. As shown in that lemma, this problem can be solved in pseudo-polynomial time via a dynamic programming approach. Therefore, the SHAP query for classification over GAMs with enumerable discrete input domains is also solvable in pseudo-polynomial time.
Appendix I Proof of Proposition 5
Proposition 5.
Let there be a NAM or an EBM over a discrete or contentious input space, then solving the CSR, and MSR queries are coNP-Complete, and the MCR query is NP-Complete.
Proof. We will divide the proof into several lemmas, beginning with NAMs and then proceeding to EBMs. For each model, we will first establish the results for CSR, followed by MCR, and finally the MSR query. We will also start with continuous domains and then explain how to extend the results to general discrete domains in each case.
Lemma 13.
Given a NAM $f$, an input $\textbf{x}\in\mathbb{F}$, and an integer $k\in\mathbb{N}$, where $f$ is defined over either a continuous input space or a general discrete input space, then obtaining the CSR query is coNP-Complete.
Proof. Membership. In the general discrete setting, the proof is straightforward. One can guess a set of $|\overline{S}|$ assignments corresponding to the features in $\overline{S}$. This is feasible because the domain of each feature $i$, denoted $\mathcal{X}_{i}$, is defined by a minimum and maximum value, each represented with at most $q$ bits. Moreover, any value in the domain can also be expressed with at most $q$ bits. Therefore, any value for a feature $i$ in $\overline{S}$ can be determined by guessing the bit values for each feature representation. Denote the guessed values for $\overline{S}$ as $\textbf{z}_{\bar{S}}$, and select arbitrary assignments for the features in $S$ (chosen from their respective domains).
This allows us to construct a vector z that includes both the guessed values for $\overline{S}$ and the arbitrarily chosen values for $S$. We then verify whether $f(\textbf{x}_{S};\textbf{z}_{\bar{S}})\neq f(\textbf{x})$, which can be done in polynomial time. If this condition holds, it implies that $S$ is not a sufficient reason for $\langle f,\textbf{x}\rangle$, thereby establishing membership in coNP.
In the continuous setting, the situation becomes more intricate, as guessing a witness assignment for the input is not feasible. This is because the size of the encoding is no longer polynomial and may not even be finite. However, as shown in a similar proof from [95], for a neural network with ReLU activations, one can, instead of guessing the input assignment, guess the activation status of each neuron (either active or inactive).
Once the activation status of each neuron is established, and assuming all remaining constraints are linear, finding a satisfying assignment for the neural network reduces to solving a linear programming problem. In our case, this indeed holds, since it involves setting lower and upper bounds for each input (corresponding to the minimum and maximum permissible values in $\mathcal{X}_{i}$ for each feature $i\in\overline{S}$) and equality constraints for features in $S$. For the output constraint, we impose a condition that it is either greater or less than zero, depending on the opposite of the classification result for $f(\mathbf{x})$. If this linear program yields a feasible solution, it demonstrates that $f(\mathbf{x}_{S};\mathbf{z}_{\bar{S}})\neq f(\textbf{x})$, hence proving membership in coNP.
Hardness. We begin by establishing hardness through the (complementary of) the neural network verification problem, which has been examined in [69] and [95], among numerous other studies. This problem, known to be NP-Complete [69, 95], is formalized as follows:
Neural Network Verification: Input: A neural network $f:\mathbb{R}^{d}\to\mathbb{R}^{c}$, a bounded continuous domain $\mathcal{X}$ over the input space, and a bounded continuous domain $\mathcal{Y}$ over the output space. Output: Yes, if there exists some $\textbf{x}\in\mathcal{X}$ and some $\mathbf{y}\in\mathcal{Y}$ such that $f(x)=y$, and No otherwise
We will prove hardness by reducing from the (complementry of) the Neural Network verification problem, which is known to be NP-Complete [69, 95]. We note that while the authors in both [69] and [95] describe this as a more general problem for wny piec-wise linear specification over the input and output (and not only for bounded domains), in practice the hardness result that were proven from the classic SAT problem were performed for specificacations of this form (bounding both the input and the outputs). Hence, the NP-Hardness of these problems holds from these reductions. Moreover, as noted by [95] — hardness remains for neural networks consisting of a single output neuron, and hence we can assume that our neural network is actually of the form $f:\mathbb{R}^{d}\to\mathbb{R}$.
We will begin by establishing an intermediate proof for our claim, which is even a stronger one that was obtained by [95]. Particularly, we will prove that the neural network verification problem remains NP-Hard even for a model with both a single input and a single output. We will later demonstrate why this hardness result also applies to the general discrete input setting.
Lemma 14.
Given a neural network $f:\mathbb{R}\to\mathbb{R}$, a bounded continuous input domain $\mathcal{X}$, and a bounded continuous output domain $\mathcal{Y}$, then solving the neural network verification problem over $f$ is NP-Hard.
We will begin by establishing hardness for the continuous domain and then demonstrate why these results extend to the general discrete domain. Now, let there be a neural network $f:\mathbb{R}^{d}\to\mathbb{R}$ with a bounded continuous input domain $\mathcal{X}$, and a bounded continuous output domain $\mathcal{Y}$. First, we will show how to reduce $f$ into a model $f^{\prime}:\mathbb{R}\to\mathbb{R}$, with some continuous bounded domain $\mathcal{X}^{\prime}$ over the (individual) input of $f^{\prime}$ and the same output domain $\mathcal{Y}$, for which it holds that there exists some $\textbf{x}\in\mathcal{X},\textbf{y}\in\mathcal{Y}$ for which $f(\textbf{x})=\textbf{y}$ if and only if there also exists some $\textbf{x}^{\prime}\in\mathcal{X^{\prime}}$ for which $f^{\prime}(\textbf{x}^{\prime})=\textbf{y}$.
We first note that given some value $o_{i}$ that is obtained by some arbitraty neuron in an MLP, it is possible to construct the calculation of $o_{j}:=|o_{i}|$ in the preceding layers by incorporating the following procedure:
$o_{j}=\text{ReLU}(o_{i})+(-1)\text{ReLU}(-o_{i})=|o_{i}|$ (24)
which can be computed by simply adding two additional neurons in the preceding layers with $1$ and 1-1 weight and connecting them both to another preceding layer. We hence, will consider, for simplicity that for any $o_{i}$, $o_{j}:=|o_{i}|$ can be computed by a constant number of preceding layers. We also note that given some $o_{i}$ we can simply compute:
$o_{j}=\text{ReLU}(o_{i})+\text{ReLU}(-o_{i})=o_{i}$ (25)
meaning that a single value of a neuron can be passed to the subsequent layers with a simple identity transformation with that construction. We will hence regard the identity construction as a valid one as well. Lastly we regard the following construction:
$o_{j}=\text{ReLU}(o_{i})-1=\text{max}(o_{i},0)$ (26)
We observe that, as previously mentioned, the proof in [95] was derived from the CNF-SAT problem, initially assuming the input space $\mathcal{X}:=\{0,1\}^{n}$. Consequently, we will specifically demonstrate hardness for the neural network verification problem involving a model of the form $f^{\prime}:\mathbb{R}\to\mathbb{R}$ by reducing it from the neural network verification problem over a model of the form $f:\{0,1\}^{n}\to\mathbb{R}$.
Furthermore, we observe that the input domain required for our model $f$ is a bounded continuous domain $\mathcal{X}$, defined by minimum and maximum values represented in binary. For simplicity and clarity, we demonstrate how to reduce a model $f$, defined over the domain $\mathcal{X}:=[0,2^{n}-1]$, to the domain $\{0,1\}^{n}$. This reduction is achieved through a binary search construction over $0,2^{n-1}]$ and can be generalized to any bounded continuous domain $\mathcal{X}$, where the minimum value of $\mathcal{X}$ corresponds to 0 and the maximum value corresponds to $2^{n-1}$.
Reducing the continuous domain $0,2^{n}-1]$ for an individual feature to the discrete domain $\{0,1\}^{n}$ for $n$ features. The construction of $f^{\prime}$ is as follows. Following the first input neuron $o^{0}$ we will connect it to a neuron in the preceding layers which calculates: $o^{1}_{0}:=\text{ReLU}(o^{0}-2^{n-1})$ and $o^{1}_{1}:=\text{ReLU}(o^{1}_{0})-\text{ReLU}(o^{1}_{0}-2^{n-1})$. We connect $o^{1}_{1}$ to the first neuron from layer $o^{\prime}_{1}$ via a linear transformation, i.e., $o^{\prime}_{1}:=o^{1}_{1}$. We also set $o^{1}_{2}:=o^{1}_{0}-\text{ReLU}(o^{1}_{1})$.
We now move on and construct another hidden layer over which we compute $o^{2}_{0}:=\text{ReLU}(o^{1}_{1}-2^{n-2})$ and $o^{2}_{1}:=\text{ReLU}(o^{1}_{1})-\text{ReLU}(o^{1}_{1}-2^{n-2})$. We connect $o^{2}_{1}$ to the second neuron in the $o^{\prime}$ layer via a linear transformation, i.e., we compute $o^{\prime}_{2}:=o^{2}_{1}$. We then again set $o^{2}_{2}:=o^{2}_{0}-\text{ReLU}(o^{2}_{1})$ and $o^{3}_{0}:=\text{ReLU}(o^{2}_{1}-2^{n-3})$, and continue performing this entire process recursively, i.e., for the $i$’th iteration in this process we construct $o^{i}_{0}:=\text{ReLU}(o^{i-1}_{2}-2^{n-i})$, $o^{i}_{1}:=\text{ReLU}(o^{i}_{1})-\text{ReLU}(o^{i}_{0}-2^{n-i})$$o^{i}_{2}:=o^{i}_{0}-\text{ReLU}(o^{i}_{1})$ and connect input $i$ in the $o^{\prime}$ layer to $o^{i}_{1}$ via a linear transformation, i.e.,: $o^{\prime i}:=o^{i}_{1}$.
The following construction takes some arbitrary assignment $o_{0}$ in the range $0,2^{n}]$ and performs a set of binary decisions, each time decreasing $2^{i-1}$ and continuing with the remainder. This process hence gives us in the first neuron $o^{\prime}_{1}$ a positive value iff $o_{0}$ is larger than $2^{n-1}$, and otherwise will be assigned a 0. The second neuron $o^{\prime}_{2}$ will get a positive value iff the remainder after decreasing $2^{n-1}$ is larger than 0, and otherwise it will be set to 0. The third neuron $o^{\prime}_{3}$ will get a positive value iff the remainder of what is left, while decreasing $2^{n-2}$ from it is larger than 0, and otherwise, it will be assigned 0. This continues for every $i$. It is straightforward to show that for a value $o^{\prime}=\lfloor x\rfloor$ for some $0\geq x\geq 2^{n}$ it holds that we can equivalentley describe $o^{\prime}$ as a binary vector of size $n$ which is equivalent to the binary value of $x$, where each “1” value in place $i$ in the binary vector corresponds to some positive assignment in place $i$ in the vector $o^{\prime}$, and each “0” value in place $i$ in the binary vector corresponds to some 0 assignment in the $o^{\prime}$ vector.
We now will describe an additional transformation over $o^{\prime}$ that we do befor connecting it to get $o^{\prime\prime}$ which we will connect to the first hidden layer of the original one in $f$ (and hence to the preceding layers in $f$ as well). First, let us observe that for the following construction:
$z:=\text{ReLU}(1-x)+\text{ReLU}(x-1)$ (27)
it holds that if $0\leq x\leq 1$ then $z=1$. We hence can define the additional hidden layer $o^{\prime\prime}$ which comes after $o^{\prime}$ such that for each $o^{\prime}_{i}$ we define:
$o^{\prime\prime}_{i}:=\text{ReLU}(1-\frac{1}{2^{n}}o^{\prime}_{i})+\text{ReLU}(\frac{1}{2^{n}}o^{\prime}_{i}-1)$ (28)
Since the range of the original individual input of $f^{\prime}$: $o^{0}$ is in $0,2^{n}]$, then we have that each remainder that is propagated into the $o^{\prime}$ layer is also in the range $0,2^{n}]$. We hence get that for any positive instance that was propagated into $o^{\prime}$ (and represents its corresponding binary vector) its output will be equal to exactly $1$ in $o^{\prime\prime}$.
We connect layer $o^{\prime\prime}$ to the remaining hidden layers from the original $f$. We note that the continuous domain $0,2^{n}]$ over the single input neuron in $o^{0}$ in $f$ is mapped exactly to any possible binary vector $\{0,1\}^{n}$ in layer $o^{\prime\prime}$. The preceding layers after $o^{\prime\prime}$ are identical between $f$ and $f^{\prime}$ and hence we have that:
$\text{max}_{x\in[0,2^{n}-1]}f^{\prime}(x)=\text{max}_{x\in\{0,1\}^{n}}f(x)\quad and\quad\text{min}_{x\in[0,2^{n}-1]}f^{\prime}(x)=\text{min}_{x\in\{0,1\}^{n}}f(x)$ (29)
We connect layer $o^{\prime\prime}$ to the remaining hidden layers from the original $f$. We note that the continuous domain $0,2^{n}]$ over the single input neuron in $o^{0}$ in $f$ is mapped exactly to any possible binary vector $\{0,1\}^{n}$ in layer $o^{\prime\prime}$. The preceding layers after $o^{\prime\prime}$ are identical between $f$ and $f^{\prime}$. Thus, for any $\textbf{x}\in[0,2^{n-1}]$, we can consider the binary representation of the lower value within the range that maps to $\{0,1\}^{n}$ (denoted as $\textbf{x}^{\prime}$), satisfying $f^{\prime}(\textbf{x}^{\prime})=f(\textbf{x})$. This demonstrates more specifically that there exists $\textbf{x}\in\mathcal{X},\textbf{y}\in\mathcal{Y}$ such that $f(\textbf{x})=\textbf{y}$ if and only if there exists some $\textbf{x}^{\prime}\in\mathcal{X}^{\prime}=\{0,1\}^{n}$ for which $f(\textbf{x})=f^{\prime}(\textbf{x}^{\prime})=\textbf{y}$. This hence concludes the proof of our Lemma.
Extending the proof to the general discrete setting. We observe that the construction we developed for $f^{\prime}$ relies on performing a binary search over the features within $0,2^{n-1}]$, mapping them into the discrete input space ${0,1}^{n}$ by associating each value with its corresponding lower value in the binary domain. If we consider the domain $\mathcal{X^{\prime}}:=[0,2^{n-1}]$ to consist solely of discrete binary values represented with $n$ bits, the mapping functions identically. In this case, instead of assigning a lower value for each mapping, the values are mapped exactly. Consequently, it is straightforward to conclude that the hardness proof remains valid when $\mathcal{X}$ is assumed to be general discrete.
Finalizing the proof. We have established that the neural network verification problem for a neural network with one input and one output, where the input domain is either continuous or general discrete, is NP-Hard. We now aim to leverage this hardness result to demonstrate that the CSR query for a NAM defined over either a general discrete or a continuous domain is coNP-Hard. We observe that the hardness proof we derived applies to a specific scenario in which the output specification $\mathcal{Y}$ mandates that the prediction of $f(\textbf{x})$ must be positive (i.e., $f(\textbf{x})\geq 0$). Consequently, we will demonstrate hardness by reducing such an instance — a neural network with a single input and output, a bounded input domain $\mathcal{X}$, where the task is to determine whether $f(\textbf{x})\geq 0$ — to the (complement of) the CSR query.
Given a model $f$ and an input specification $\mathcal{X}$, we construct a NAM $f^{\prime}$ consisting of only one MLP, $f_{1}:=f$. We start by selecting an arbitrary assignment $\textbf{x}\in\mathcal{X}$ and checking if $f(\textbf{x})<0$. If this condition is satisfied, we define the weight of the corresponding model as $\beta_{1}:=1$ and set the bias term as $\beta_{0}:=0$. Furthermore, we initialize the subset $S:=\emptyset$. It is easy to see that if there exists a satisfying assignment $\textbf{z}\in\mathcal{X}$ for which $f(\textbf{z})\geq 0$, then $n]:=\{1\}$ is a contrastive reason, since $1=f^{\prime}(\textbf{z})\neq f^{\prime}(\textbf{x})=0$ and thus $\emptyset$ is not a sufficient reason for $\langle f^{\prime},\textbf{x}\rangle$. In summary, we have constructed an instance of the CSR query where $S$ (the empty set) acts as a sufficient reason for $\langle f^{\prime},\textbf{x}\rangle$ if and only if no satisfying assignment exists for the neural network $f$.
For the second case, where $f(\textbf{x})\geq 0$, we construct $f^{\prime}$ in the same way, but this time setting the weight $\beta_{1}:=-1$, while keeping $\beta_{0}:=0$ and $S:=\emptyset$. Similarly, we observe that if there exists a satisfying assignment $\textbf{z}\in\mathcal{X}$ such that $f(\textbf{z})\geq 0$, then $n]:=\{1\}$ serves as a contrastive reason, as $0=f^{\prime}(\textbf{z})\neq f^{\prime}(\textbf{x})=1$, implying that $\emptyset$ is not a sufficient reason for $\langle f^{\prime},\textbf{x}\rangle$. To summarize, we have constructed a CSR query instance where $S$ (the empty set) serves as a sufficient reason for $\langle f^{\prime},\textbf{x}\rangle$ if and only if no satisfying assignment exists for the neural network $f$.
This establishes that obtaining the CSR query for a NAM, where $\mathcal{X}$ is defined as either a general discrete or a continuous input domain, is coNP-Hard.
Lemma 15.
Given a NAM $f$, an input $\textbf{x}\in\mathbb{F}$, and an integer $k\in\mathbb{N}$, where $f$ is defined over either a continuous input space or a general discrete input space, then the MCR query is NP-Complete.
Proof. Membership. For proving membership in NP, we follow the exact same procedure used for proving membership in coNP for the CSR query in Lemma 13. However, instead of verifying that $f(\textbf{x})\neq f(\textbf{x}_{S};\textbf{z}_{\bar{S}})$, we check that $f(\textbf{x})=f(\textbf{x}_{S};\textbf{z}_{\bar{S}})$ while introducing an additional constraint that $|S|\leq k$. Since these conditions can still be determined in polynomial time (with the difference being that we now demonstrate the existence of an instance, rather than the lack of one), this establishes membership in NP.
Hardness. We have proven in Lemma 14 that the neural network verification problem for a neural network with one input and one output, where the input domain is either continuous or general discrete, is NP-Hard. We now aim to leverage this hardness result to demonstrate that the MCR query for a NAM defined over either a general discrete or a continuous domain is NP-Hard.
We observe that the hardness proof we derived applies to a specific scenario in which the output specification $\mathcal{Y}$ mandates that the prediction of $f(\textbf{x})$ must be positive (i.e., $f(\textbf{x})\geq 0$). Consequently, we will demonstrate hardness by reducing such an instance—a neural network with a single input and output, a bounded input domain $\mathcal{X}$, where the task is to determine whether $f(\textbf{x})\geq 0$ — to the MCR query.
Given a model $f$ and an input specification $\mathcal{X}$, we construct a NAM $f^{\prime}$ consisting of only one MLP, $f_{1}:=f$. We start by selecting an arbitrary assignment $\textbf{x}\in\mathcal{X}$ and checking if $f(\textbf{x})<0$. If this condition is satisfied, we define the weight of the corresponding model as $\beta_{1}:=1$ and set the bias term as $\beta_{0}:=0$. Furthermore, we set $k:=1$. It is easy to see that if there exists a satisfying assignment $\textbf{z}\in\mathcal{X}$ for which $f(\textbf{z})\geq 0$, then $n]:=\{1\}$ is a contrastive reason, since $1=f^{\prime}(\textbf{z})\neq f^{\prime}(\textbf{x})=0$ and thus there exists a contrastive reason of size $1$ for $\langle f^{\prime},\textbf{x}\rangle$.
For the second case, where $f(\textbf{x})\geq 0$, we construct $f^{\prime}$ in the same way, but this time setting the weight $\beta_{1}:=-1$, while keeping $\beta_{0}:=0$ and, again, setting $k:=1$. Similarly, we observe that if there exists a satisfying assignment $\textbf{z}\in\mathcal{X}$ such that $f(\textbf{z})\geq 0$, then $n]:=\{1\}$ serves as a contrastive reason, as $0=f^{\prime}(\textbf{z})\neq f^{\prime}(\textbf{x})=1$, implying that there exists a contrastive reason of size $1$. To summarize, we have constructed an MCR query instance where there exists a contrastive reason of size $k:=1$ if and only if there exists a satisfying assignment for the neural network $f$.
This establishes that obtaining the MCR query for a NAM, where $\mathcal{X}$ is defined as either a general discrete or a continuous input domain, is NP-Hard.
Lemma 16.
Given a NAM $f$, an input $\textbf{x}\in\mathbb{F}$, and an integer $k\in\mathbb{N}$, where $f$ is defined over either a continuous input space or a general discrete input space, then obtaining the MSR query is coNP-Complete.
Proof. Hardness. We have proven in Lemma 14 that the neural network verification problem for a neural network with one input and one output, where the input domain is either continuous or general discrete, is NP-Hard. We can now leverage this hardness result to demonstrate that the MSR query for a NAM defined over either a general discrete or a continuous domain is coNP-Hard.
We, again, can observe that the hardness proof we derived applies to a specific scenario in which the output specification $\mathcal{Y}$ mandates that the prediction of $f(\textbf{x})$ must be positive (i.e., $f(\textbf{x})\geq 0$). Consequently, we will demonstrate hardness by reducing such an instance — a neural network with a single input and output, a bounded input domain $\mathcal{X}$, where the task is to determine whether $f(\textbf{x})\geq 0$ — to the (complement of) the MSR query.
Given a model $f$ and an input specification $\mathcal{X}$, we construct a NAM $f^{\prime}$ consisting of only one MLP, $f_{1}:=f$. We start by selecting an arbitrary assignment $\textbf{x}\in\mathcal{X}$ and checking if $f(\textbf{x})<0$. If this condition is satisfied, we define the weight of the corresponding model as $\beta_{1}:=1$ and set the bias term as $\beta_{0}:=0$. Furthermore, we set $k:=0$. It is easy to see that if there exists a satisfying assignment $\textbf{z}\in\mathcal{X}$ for which $f(\textbf{z})\geq 0$, then $n]:=\{1\}$ is a contrastive reason, since $1=f^{\prime}(\textbf{z})\neq f^{\prime}(\textbf{x})=0$ and thus there does not exist a sufficient reason of size 0 for $\langle f^{\prime},\textbf{x}\rangle$.
For the second case, where $f(\textbf{x})\geq 0$, we construct $f^{\prime}$ in the same way, but this time setting the weight $\beta_{1}:=-1$, while keeping $\beta_{0}:=0$ and, again setting $k:=0$. Similarly, we observe that if there exists a satisfying assignment $\textbf{z}\in\mathcal{X}$ such that $f(\textbf{z})\geq 0$, then $n]:=\{1\}$ serves as a contrastive reason, as $0=f^{\prime}(\textbf{z})\neq f^{\prime}(\textbf{x})=1$, implying that there exists a contrastive reason of size $1$, and hence there does not exist a sufficient reason of size 0. To summarize, we have constructed an MSR query instance where there does not exist a sufficient reason of size $k:=0$ if and only if there exists a satisfying assignment for the neural network $f$.
This establishes that obtaining the MSR query for a NAM, where $\mathcal{X}$ is defined as either a general discrete or a continuous input domain, is coNP-Hard. We note that a recent result by [22] strengthens this bound for the optimization variant, showing that the continuous version of the problem is in fact OptP[$\mathcal{O}(\log n)$]-hard under metric reductions (see Section B.5 of the appendix).
Regarding membership, although showing that the problem lies in $\Sigma_{2}^{P}$ is straightforward (as the same task is known to be in $\Sigma_{2}^{P}$ for general neural networks [15]), we refer the reader to a recent result in [22], which shows that the optimization variant of the problem is in $\mathrm{FP}^{\mathrm{NP}}[\mathcal{O}(n)]$ (see Section B.5 of the appendix). This, in turn, implies that the corresponding decision variant considered here belongs to $\Delta_{2}^{P}$.
Having completed the proofs for all NAM instances, we now turn to proving the complexity results for EBMs. We will break down the proof into separate lemmas, each handling the complexity argument for one of the explanation types (CSR, NCR, and MSR), starting with the continuous case and then extending to the general discrete case.
Lemma 17.
Given an EBM over either a continuous or general discrete input space, then the CSR query is coNP-Complete.
Proof. Membership. Although we can directly prove membership for EBMs, we choose to establish it via a reduction to NAMs. Given that the same task is already shown to be coNP-complete for NAMs, and since coNP is closed under polynomial-time reductions, this approach immediately yields membership in coNP. A key advantage of this proof is that it extends naturally to other queries — such as MCR and MSR — since the same reduction applies to them as well.
We can directly reduce any EBM to a NAM by transforming each boosted tree ensemble $f_{i}$ into a neural network $f^{\prime}_{i}$. This is done by converting each decision tree into a neural network, following the method proposed in [15, 17]. To account for the ensemble structure, we introduce an additional hidden layer that assigns a weight $\phi_{j}$ to each tree $j$. A standard weighted sum over the outputs then yields a neural network equivalent to the original boosted ensemble. This transformation preserves equivalence for any given input.
Hardness. We perform a reduction from the classic tautology problem (TAUT) for 3-DNFs (which is coNP-Complete), defined as follows:
TAUT (Tautology): Input: A boolean 3-DNF formula $\psi:=t_{1}\vee t_{2}\vee\ldots t_{m}$. Output: Yes, if $\psi$ is a tautology and No otherwise.
We will begin by proving the hardness for the general discrete input space and then explain how to extend the results to the continuous input space. Given a DNF $\phi:=t_{1}\vee t_{2}\vee\ldots\vee t_{m}$, we will construct an EBM $f:=\langle f_{1}\rangle$, i.e., the EBM $f$ is composed of only one boosted tree. It is worth noting that reducing DNFs to boosted trees has been proposed in previous work [63]. However, our situation differs because the boosted tree $f_{1}$ is defined over only one input feature $x_{1}$ and operates on some input space $\mathcal{X}$, i.e., $f_{1}:\mathcal{X}\to\mathbb{R}$ and $f:\mathcal{X}\to\{0,1\}$.
We recall that $\phi$ is defined over a set of $n$ literals $X_{1},X_{2},\ldots,X_{n}$ and $m$ clauses. Given $\phi$, we will construct a boosted tree model $f_{1}$ (which will thereby define $f:=\langle f_{1}\rangle$) and set the general discrete input space $\mathcal{X}_{i}$ to encompass any possible integer within the range $0,2^{n}-1]$. Each feature $X_{i}$ will be associated with its corresponding bit representation in a vector containing $n$ bits.
We note that $f_{1}$ has one input, $\textbf{x}_{1}$, defined over the domain $\mathcal{X}_{i}$. The model $f_{1}$ will be an ensemble comprising $m$ decision trees. Each decision tree’s input is the feature $\textbf{x}_{1}$. Consider a clause $t_{i}:=c_{1}\wedge c_{2}\wedge c_{3}$, where each conjunct is a clause within $X_{1},\ldots,X_{n}$ or its negation. For each conjunct $t_{i}:=c_{1}\wedge c_{2}\wedge c_{3}$, we will construct a decision tree corresponding to that conjunct.
This decision tree will have three splits, one for each feature $c_{1},c_{2},c_{3}$. Assume, without loss of generality, that these features are represented by $X_{j}$, $X_{k}$, and $X_{l}$, where each feature may also be represented negatively as $\bar{X_{j}}$, $\bar{X_{k}}$, or $\bar{X_{l}}$. The splits of the tree will check conditions such as $\textbf{x}_{1}\geq 2^{j}$, $\textbf{x}_{1}\geq 2^{k}$, or $\textbf{x}_{1}\geq 2^{l}$ (representing literals $X_{j}$, $X_{k}$, $X_{l}$, respectively), or $\textbf{x}_{1}<2^{j}$, $\textbf{x}_{1}<2^{k}$, or $\textbf{x}_{1}<2^{l}$ (representing literals $\bar{X_{j}}$, $\bar{X_{k}}$, $\bar{X_{l}}$, respectively). If all three conditions are satisfied, the tree will output a value of $1$; otherwise, it will output 0.
For example, consider the conjunct $t_{5}:=X_{j}\wedge\bar{X_{k}}\wedge X_{l}$. In this scenario, tree number $5$ in the constructed boosted tree ensemble $f_{1}$ will perform three splits: first, it will check if $\textbf{x}_{i}\geq 2^{j}$, then whether $\textbf{x}_{i}<2^{j}$, and finally whether $\textbf{x}_{i}\geq 2^{l}$. If all three conditions hold true, the tree will output a value of $1$. Otherwise, it will output 0. The $\alpha$ terms (the weights of each tree in $f_{1}$) will all be set to $1$. For the intercept terms of $f$, we will set $\beta_{1}=1$ and $\beta_{0}=-1$. This defines the full construction of the EBM $f$. Additionally, we will set $S:=\emptyset$. Finally, we construct the input x, which is any integer within the range $0,2^{n}-1]$, as follows: First, we form a binary vector $\textbf{x}^{\prime}\in\{0,1\}^{n}$ in the following way: we start by choosing an arbitrary conjunct from $c_{1},\ldots,c_{m}$. For a specific conjunct, we assign the corresponding literals to their respective bits in the vector, while assigning any values to the remaining bits. For instance, consider the clause $c_{i}:=X_{1}\wedge\bar{X_{5}}\wedge X_{8}$. In this case, we create a vector $\textbf{x}^{\prime}$ where $\textbf{x}^{\prime}_{1}:=1$, $\textbf{x}^{\prime}_{5}:=0$, and $\textbf{x}^{\prime}_{8}:=1$, while the other features in $\textbf{x}^{\prime}$ (besides features 1, 5, and 8) are assigned arbitrary values. We then define the input x as the integer represented by the binary vector $\textbf{x}^{\prime}$. In summary, the entire reduction constructs $\langle f,\textbf{x},S\rangle$, where $f$ is the constructed EBM defined over the input space $\mathcal{X}_{i}$ for the single feature $\textbf{x}_{i}$ of $f$, x is the constructed input, and $S:=\emptyset$ is the constructed (empty) subset.
We will now prove that $\phi$ is a tautology if and only if $S:=\emptyset$ is a sufficient reason for $\langle f,\textbf{x}\rangle$. From our construction of x (the integer representation of the binary vector $\textbf{x}^{\prime}$, which expands binary bits representing literals from a given conjunct), it follows that for the specific conjunct $c_{i}$ over which $\textbf{x}^{\prime}$ was constructed, the corresponding tree in the ensemble will reach a terminal node with value 1. Consequently, since all other trees output either 0 or 1, and given the intercept term $\beta_{0}=-1$, we have $f(\textbf{x})=1$. Knowing that $f(\textbf{x})=1$, proving that $S=\emptyset$ is a sufficient reason is equivalent to showing that for any z, it holds that $f(\textbf{x})=1=f(\textbf{z})$. Conversely, proving that $S=\emptyset$ is not a sufficient reason amounts to demonstrating the existence of some z such that $f(\textbf{x})=1\neq f(\textbf{z})=0$.
From our construction, we know that for any assignment of literals $\textbf{z}^{\prime}\in\{0,1\}^{n}$ over $\phi$, if a conjunct $c_{i}$ is satisfied, propagating z (the integer representation of the binary vector $\textbf{z}^{\prime}$) through the $i$-th tree of the ensemble will yield a terminal node with value 1. Since the outputs of all other trees are either 0 or 1, it follows that $f_{1}(\textbf{z})\geq 1$. Including the intercept $\beta_{0}=-1$, we get $f(\textbf{z})=1=f(\textbf{x})$.
On the other hand, if for some assignment $\textbf{z}\in\{0,1\}^{n}$, no conjunct in $\phi$ is satisfied, all trees will reach terminal nodes with value 0, resulting in $f(\textbf{z})=0\neq 1=f(\textbf{x})$. Thus, if there exists an assignment over $\phi$ that evaluates to false (i.e., $\phi$ is not a tautology), there will exist an assignment z such that $f(\textbf{z})=0\neq 1=f(\textbf{x})$, implying that $S:=\emptyset$ is not a sufficient reason for $\langle f,\textbf{x}\rangle$. Conversely, if no such assignment exists (i.e., $\phi$ is a tautology), then for any z, it holds that $f(\textbf{x})=1=f(\textbf{z})$, implying that $\emptyset$ is a sufficient reason. This concludes the proof.
Finally, we note that, for simplicity, we assumed $\mathcal{X}_{1}$, the domain of the single input $\textbf{x}_{1}$, to be a general discrete domain containing all integer values within $0,2^{n}-1]$. However, the result also holds if $\mathcal{X}_{1}$ is defined as the entire continuous domain $0,2^{n}-1]$. Notably, in our construction of the single tree ensemble $f_{1}$ within the EBM, the decision rule checking whether $\textbf{x}\geq 2^{k}$ or $\textbf{x}<2^{k}$ applies equally to the entire continuous domain. Thus, the hardness results extend to the continuous domain as well.
Lemma 18.
Given an EBM over either a continuous or general discrete input space, then obtaining the MCR query is NP-Complete.
Proof. For membership, we can use the same reasoning as in Lemma 17, where the given EBM is reduced to an equivalent NAM. Since NP is closed under polynomial-time reductions, Lemma 18 establishes membership in NP.
For hardness, we reduce from the (complement) of the CSR query for EBMs, which we have shown to be coNP-Complete. Specifically, we established coNP-hardness for the CSR query by reducing from the TAUT problem to CSR when $S:=\emptyset$. Consequently, we prove hardness by reducing from CSR in the special case where $S:=\emptyset$. Given an instance $\langle f,\textbf{x},S:=\emptyset\rangle$, where $f$ is an EBM, we construct an instance $\langle f,\textbf{x},k=n\rangle$, where $f$ remains the same EBM, and $k=n$ corresponds to setting $k$ of the MCR query to match the input dimension size. We observe that $S=\emptyset$ is a sufficient reason for $\langle f,\textbf{x}\rangle$ if and only if $\overline{S}$ is not a contrastive reason for $\langle f,\textbf{x}\rangle$. Thus, $S=\emptyset$ is a sufficient reason for $\langle f,\textbf{x}\rangle$ if and only if no contrastive reason of size $n$ exists for $\langle f,\textbf{x}\rangle$, thereby establishing the reduction.
Lemma 19.
Given an EBM over either a continuous or general discrete input space, then the MSR query is coNP-Complete.
Proof. For membership, the same reasoning as in Lemma 17 and Lemma 18 applies, where the given EBM is reduced to an equivalent NAM. As coNP is closed under polynomial-time reductions, Lemma 17 confirms membership in coNP.
For hardness, we reduce from the CSR query for EBMs, which has been shown to be coNP-Complete. Specifically, coNP-hardness for the CSR query was established by reducing from the TAUT problem to CSR when $S:=\emptyset$. Accordingly, hardness is proved by reducing from CSR in the specific case where $S:=\emptyset$. For an instance $\langle f,\textbf{x},S:=\emptyset\rangle$, where $f$ is an EBM, we construct an instance $\langle f,\textbf{x},k=0\rangle$, where $f$ remains the same EBM and $k=0$ corresponds to setting $k$ of the MSR query. It is observed that $S=\emptyset$ if and only if a sufficient reason of size 0 exists for $\langle f,\textbf{x}\rangle$, thereby establishing the reduction.
Appendix J Proof of Proposition 6
Proposition 6.
Given a NAM, a Smooth GAM or an EBM under an enumerable discrete or general discrete setting, the FR query is coNP-Complete.
Proof. Membership. Given a GAM $f$ defined over either an enumerable discrete or general discrete input space, and an input feature $i\in[n]$, we can guess two input assignments for feature $i$: $\textbf{x}_{i},\textbf{z}_{i}$ such that $\textbf{x}_{i}\neq\textbf{z}_{i}$. This is achievable due to the discrete nature of the input space: by selecting a single input for each coordinate in the enumerable discrete case, or by choosing a binary vector in the general discrete case. We can additionally guess some assignment for the features $n]\setminus\{i\}$, which we will denote by $\textbf{x}^{\prime}$ (and the correctness holds for the same reasoning). Now, if $f(\textbf{x}^{\prime}_{[n]\setminus\{i\}};\textbf{x}_{i})\neq f(\textbf{x}^{\prime}_{[n]\setminus\{i\}};\textbf{z}_{i})$, then by definition, $i$ is not redundant with respect to $\langle f,i\rangle$, thereby completing the membership proof. We note that this membership result holds regardless of the specific type of GAM — it applies to any GAM, provided that the inference of its components can be performed in polynomial time. Therefore, it directly extends to models such as NAMs, EBMs, and Smooth GAMs.
Hardness. The study in [17] established coNP-Hardness for the FR query (referred to there as G-FR) for linear classifiers over a binary input space. Given that this represents a specific instance of the enumerable discrete input setting, where the explicit values $\mathcal{X}_{i}$ for each feature $\textbf{x}_{i}$ are $\{0,1\}$ (and is therefore also a particular case of the general discrete input setting), the coNP-hardness results apply to these settings as well. We note that linear models are a particular instance of Smooth GAMs, NAMs and EBMs, where each component $f_{i}$ is taken to be the identity function. Since each of these model types (neural networks, tree ensembles, and piecewise polynomial splines) can trivially be represented as constant identity functions that always produce the same output, the hardness result naturally extends to EBMs, Smooth GAMs, and NAMs.
Appendix K Proof of Proposition 7
Proposition 7.
Given NAMs and EBMs over discrete or continuous settings, computing SHAP for regression tasks is #\#P-Complete. Moreover, given any GAM over an enumerable discrete, discrete or continuous setting - computing the CC and SHAP queries for classification are #\#P-Complete.
Proof. We will divide the upcoming proof into two distinct lemmas.
Lemma 20.
Given NAMs and EBMs over discrete or continuous settings, computing SHAP for regression tasks is #\#P-Hard.
Proof. We will start by proving the results for NAMs. First, similarly to the proof in Lemma 5, we will assume we are dealing with a single input-output neural network, and hence, hardness for this scenario will directly apply to full NAMs. We base the reduction based on the results obtained by [8] and [102] which showed that computing SHAP for a model $f$ and an input x is at least as hard as the model counting problem for $f$, under the uniform distribution assumption (and this clearly shows hardness for the more general distribution setting). CNF formulas can be reduced in polynomial time to neural networks, as was shown by multiple studies [15, 69, 96, 17] (and this claim actually holds even for boolean circuits [15]). As shown by [69] and later refined by [96], a neural network defined over a continuous domain can be reduced to the discrete case, thereby transferring the hardness results to the continuous setting as well.
In Lemma 5, we showed that a neural network verification problem for a full network can be reduced to one involving a single input-output pair. Since the former is known to be hard via a reduction from CNF-SAT [69, 96], this implies that the counting version of CNF-SAT can likewise be reduced to counting the number of satisfying assignments in a single input-output neural network. This yields a parsimonious reduction, establishing #\#P-hardness. As EBMs can be reduced from NAMs (as we show in Lemma 5), this hardness result extends to EBMs as well.
Lemma 21.
Given any GAM over an enumerable discrete, discrete or continuous setting - computing the CC and SHAP queries for classification are #\#P-Hard.
Proof. Hardness results for GAMs persist because they apply even to linear classification models, which are a special case of GAMs. Consequently, the hardness remains for GAMs. Specifically, for the CC query, this result was established by [15], and for the SHAP query, it was demonstrated by [102]. Linear models can be viewed as a special case of Smooth GAMs, NAMs, and EBMs, where each component function $f_{i}$ is simply the identity function. Since neural networks, tree ensembles, and piecewise polynomial splines can trivially replicate this behavior by acting as constant identity functions that yield a fixed output, the hardness result directly extends to EBMs, Smooth GAMs, and NAMs.